### v5

# Concentration of discrepancy-based approximate Bayesian computation via Rademacher complexity

Abstract
There has been an increasing interest on summary-free solutions for approximate Bayesian computation (abc) which replace distances among summaries with discrepancies between the empirical distributions of the observed data and the synthetic samples generated under the proposed parameter values. The success of these strategies has motivated theoretical studies on the limiting properties of the induced posteriors. However, there is still the lack of a theoretical framework for summary-free abc that (i) is unified, instead of discrepancy-specific, (ii) does not necessarily require to constrain the analysis to data generating processes and statistical models meeting specific regularity conditions, but rather facilitates the derivation of limiting properties that hold uniformly, and (iii) relies on verifiable assumptions that provide more explicit concentration bounds clarifying which factors govern the limiting behavior of the abc posterior. We address this gap via a novel theoretical framework that introduces the concept of Rademacher complexity in the analysis of the limiting properties for discrepancy-based abc posteriors, including in non-i.i.d. and misspecified settings. This yields a unified theory that relies on constructive arguments and provides more informative asymptotic results and uniform concentration bounds, even in settings not covered by current studies. These advancements are obtained by relating the asymptotic properties of summary-free abc posteriors to the behavior of the Rademacher complexity associated with the chosen discrepancy within the family of integral probability semimetrics (ips). The ips class extends summary-based distances, and also includes the widely-implemented Wasserstein distance and maximum mean discrepancy (mmd), among others. As clarified in specialized theoretical analyses of popular ips discrepancies and via illustrative simulations, this new perspective improves the understanding of summary-free abc.
62F15,
62C10,
62E20,
ABC,
Integral probability semimetrics,
MMD,
Rademacher complexity,
Wasserstein distance,
keywords:
[class=MSC2020]
keywords:
[…]
, and
1 Introduction
The growing complexity of statistical models in modern applications not only yields intractable likelihoods, but also raises substantial challenges in the identification of effective summary statistics (see e.g., Fearnhead and Prangle, 2012; Marin et al., 2014; Frazier et al., 2018). Such drawbacks have motivated an increasing adoption of abc solutions, along with a shift away from summary-based implementations (see e.g., Marin et al., 2012) and towards summary-free strategies relying on discrepancies among the empirical distributions of the observed and synthetic data (see e.g., Drovandi and Frazier, 2022). These solutions provide samples from an approximate posterior distribution for the parameter of interest ${\theta\in\Theta\subseteq\mathbb{R}^{p}}$ under the only requirement that simulating from the assumed model $\mu_{\theta}$ is feasible. This is achieved by retaining all those values of $\theta$, drawn from the prior, that produced synthetic samples $z_{1:m}=(z_{1},\ldots,z_{m})$ from $\mu_{\theta}$ whose empirical distribution is close enough to the one of the observed data $y_{1:n}=(y_{1},\ldots,y_{n})$, under the chosen discrepancy.
Remarkable examples of the above implementations are abc versions that employ maximum mean discrepancy (mmd) (Park, Jitkrittum and Sejdinovic, 2016), Kullback–Leibler (kl) divergence (Jiang, Wu and Wong, 2018), Wasserstein distance (Bernton et al., 2019), energy statistic (Nguyen et al., 2020), Hellinger and Cramer–von Mises distances (Frazier, 2020), and $\gamma$-divergence (Fujisawa et al., 2021); see also Gutmann et al. (2018), Forbes et al. (2021) and Wang, Kaji and Rockova (2022) for additional examples of summary-free abc strategies. By overcoming the need to pre-select summaries, all these solutions reduce the potential information loss of summary-based abc, thereby yielding improved performance in simulation studies and illustrative applications. These promising empirical results have motivated active research on the theoretical properties of the induced abc posterior, with a main focus on the limiting behavior under different asymptotic regimes for the tolerance threshold and the sample size (Jiang, Wu and Wong, 2018; Bernton et al., 2019; Nguyen et al., 2020; Frazier, 2020; Fujisawa et al., 2021). Among these regimes, of particular interest are the two situations in which the abc threshold is either fixed or progressively shrinks as both $n$ and $m$ diverge. In the former case, the theory focus is on assessing whether the control on the discrepancy among the empirical distributions established by the selected abc threshold yields a pseudo-posterior whose asymptotic form guarantees the same threshold-control on the discrepancy among the underlying truths (e.g., Jiang, Wu and Wong, 2018). The latter addresses instead the more challenging theoretical question on whether a suitably-decaying abc threshold can provide meaningful rates of concentration for the sequence of abc posteriors around those $\theta$ values yielding a $\mu_{\theta}$ close enough to the data generating process $\mu^{*}$, as $n$ and $m$ diverge (e.g., Bernton et al., 2019).
Available contributions along these lines of research have the merit of providing theoretical support to several versions of summary-free abc. However, current theory is often tailored to the specific discrepancy analyzed, and generally relies on difficult-to-verify existence assumptions and concentration inequalities that constrain the analysis, either implicitly or explicitly, to data generating processes and statistical models which meet suitable regularity conditions, thereby lacking results that hold uniformly. Recalling Bernton et al. (2019) and Nguyen et al. (2020), this also yields concentration bounds involving sequences of control functions which are not made explicit. As such, although convergence and concentration can still be proved, the core factors that govern these asymptotic properties remain yet unexplored, thus limiting the methodological impact of current theory, while hindering the derivation of novel informative results in more challenging settings. For example, the available theoretical studies pre-assume that the discrepancy among the empirical distributions $\hat{\mu}_{z_{1:m}}$ and $\hat{\mu}_{y_{1:n}}$ suitably converges to the one among the corresponding truths $\mu_{\theta}$ and $\mu^{*}$, or, alternatively, that both $\hat{\mu}_{z_{1:m}}$ and $\hat{\mu}_{y_{1:n}}$ converge, under the selected discrepancy, to $\mu_{\theta}$ and $\mu^{*}$, respectively. While these assumptions can be verified under suitable conditions and for specific discrepancies, as clarified within Sections 2 and 3, an in-depth theoretical understanding of summary-free abc necessarily requires relating convergence to the learning properties of the selected discrepancy, rather than pre-assuming it. Such a more precise but yet-unexplored theoretical treatment has the potentials to shed light on the factors that govern the limiting behavior of discrepancy-based abc posteriors. In addition, it could possibly provide more general, verifiable and explicit sufficient conditions under which popular discrepancies are guaranteed to ensure that the previously pre-assumed convergence and concentration hold uniformly over models and data generating processes, while facilitating the study of the limiting behavior of discrepancy-based abc posteriors in more general situations where these convergence guarantees may lack.
In this article, we address the above gap by introducing an innovative theoretical framework which analyses the limiting properties of discrepancy-based abc posteriors, under a unified perspective and for different asymptotic regimes, through the concept of Rademacher complexity (see e.g., Wainwright, 2019, Chapter 4), within the general class of integral probability semimetrics (ips) (e.g., Müller, 1997; Sriperumbudur et al., 2012). Such a class naturally generalizes distances among summaries and includes the widely-implemented mmd and Wasserstein distance, among others. As clarified in Sections 23 and in Appendix C of the Supplementary Material, this perspective, which to the best of our knowledge is novel within abc, allows to derive unified, informative and uniform concentration bounds for abc posteriors under several discrepancies, in possibly misspecified and non-i.i.d. contexts. Moreover, it relies on more constructive arguments that clarify under which sufficient conditions a discrepancy within the ips class guarantees uniform convergence and concentration of the induced abc posterior; i.e., without necessarily requiring suitable regularity conditions for the underlying data generating process $\mu^{*}$ and the assumed statistical model. This yields an important theoretical and methodological advancement, since $\mu^{*}$ is often unknown in practice and, hence, verifying regularity conditions on the data generating process is generally unfeasible.
Crucially, the theoretical framework we introduce allows to prove informative theoretical results even in yet-unexplored settings that possibly lack those convergence guarantees assumed in the literature. More specifically, in these settings we derive novel upper and lower bounds for the limiting acceptance probabilities that clarify in which contexts the abc posterior is still well-defined for $n$ large enough. When this is the case, it is further possible to obtain informative supersets for the support of such a posterior. These show that, when relaxing standard convergence assumptions employed in state-of-the-art theoretical studies, the control established by a fixed abc threshold on the discrepancy among the empirical distributions does not necessarily translate, asymptotically, into the same control on the discrepancy among the corresponding truths, but rather yields an upper bound equal to the sum between the abc threshold and a multiple of the Rademacher complexity, namely a measure of richness of the class of functions that uniquely identify the chosen ips; see Section 3.1.
The above results clarify the fundamental relation among the limiting behavior of abc posteriors and the learning properties of the chosen discrepancy, when measured via Rademacher complexity. In addition, the bounds derived clarify that a sufficient condition to recover a limiting pseudo-posterior with the same threshold-control on the discrepancy among the truths as the one enforced on the corresponding empirical distributions, is that the selected discrepancy has a Rademacher complexity vanishing to zero in the large-data limit. As proved within Section 3.2, this setting also allows constructive derivations of novel, informative and uniform concentration bounds for discrepancy-based abc posteriors in the challenging regime where the threshold shrinks towards zero as both $m$ and $n$ diverge. This is facilitated by the existence of meaningful upper bounds for the Rademacher complexity associated to popular abc discrepancies, along with the availability of constructive conditions for the derivation of these bounds (Sriperumbudur et al., 2012). Such results leverage fundamental connections among the Rademacher complexity and other key quantities in statistical learning theory, such as the Vapnik–Chervonenkis (vc) dimension and the notion of uniform Glivenko–Cantelli classes (e.g., Wainwright, 2019, Chapter 4). This yields an improved understanding of the factors that govern the concentration of discrepancy-based abc posteriors under a unified perspective that further allows to (i) quantify rates of concentration and (ii) directly translate any advancement on Rademacher complexity into novel abc theory. Section 4 illustrates point (i) through a specific focus on mmd with routinely-implemented bounded kernels, and also clarifies that in the absence of guarantees on uniformly-vanishing Rademacher complexity (e.g., under Wasserstein distance in unbounded data spaces) concentration results can be still derived, but at the expense of regularity conditions for the data generating process $\mu^{*}$ and the assumed model. Point (ii) is clarified in Appendix C of the Supplementary Material, where we extend the theory from Section 3 to non-i.i.d. settings, leveraging results in Mohri and Rostamizadeh (2008) on the Rademacher complexity for $\beta$-mixing processes (e.g., Doukhan, 1994).
The illustrative simulation studies in Section 5 show that the theoretical results derived in Sections 34 find empirical evidence in practice, including also in scenarios characterized by model misspecification and data contamination (theoretical and empirical results under non-i.i.d. data generating processes can be found in Appendix C of the Supplementary Material). These findings suggest that discrepancies with guarantees of uniformly-vanishing Rademacher complexity provide a robust and sensible choice when the assumed statistical model and/or the underlying data generating process do not necessarily meet specific regularity conditions, or it is not possible to verify these conditions. This is a common situation in applications, since the data generating process is unknown in practice.
As discussed in Section 6, the unexplored bridge between discrepancy-based abc and the Rademacher complexity introduced in this article can also be leveraged to derive even more general theory by exploiting the active literature on the Rademacher complexity. For example, combining our perspective with the recent unified treatment of ips and $f$-divergences (Agrawal and Horel, 2021; Birrell et al., 2022) might set the premises to derive similarly-interpretable and general results for other discrepancies employed within abc, such as the Kullback–Leibler divergence (Jiang, Wu and Wong, 2018) and Hellinger distance (Frazier, 2020). More generally, our contribution can have implications even beyond abc, in particular on generalized Bayesian inference via pseudo-posteriors based on discrepancies (Bissiri, Holmes and Walker, 2016; Chérief-Abdellatif and Alquier, 2020; Matsubara et al., 2022; Frazier, Knoblauch and Drovandi, 2024). Proofs and additional results can be found in the Supplementary Material.
2 Integral probability semimetrics and Rademacher complexity
Denote with $y_{1:n}=(y_{1},\ldots,y_{n})\in\mathcal{Y}^{n}$ an i.i.d. sample from $\mu^{*}\in\mathcal{P}(\mathcal{Y})$, where $\mathcal{P}(\mathcal{Y})$ is the space of probability measures on $\mathcal{Y}$, and assume that $\mathcal{Y}$ is a metric space endowed with distance $\rho$; see Appendix C in the Supplementary Material for extensions to non-i.i.d. settings. Given a statistical model $\{\mu_{\theta}:\theta\in\Theta\subseteq\mathbb{R}^{p}\}$ in $\mathcal{P}(\mathcal{Y})$ and a prior distribution $\pi$ on $\theta$, rejection abc iteratively samples $\theta$ from $\pi$, draws a synthetic i.i.d. sample $z_{1:m}=(z_{1},\ldots,z_{m})$ from $\mu_{\theta}$, and retains $\theta$ as a sample from the abc posterior if the discrepancy $\Delta(z_{1:m},y_{1:n})$ between $z_{1:m}$ and $y_{1:n}$ is below a user-selected threshold $\varepsilon_{n}\geq 0$. Although $m$ may differ from $n$, here we follow common practice in abc theory (e.g., Bernton et al., 2019; Frazier, 2020) and set ${m=n}$. Rejection abc does not sample from the exact posterior $\pi_{n}(\theta)\propto\pi(\theta)\mu_{\theta}^{n}(y_{1:n}),$ but rather from the abc posterior defined as
$\pi_{n}^{(\varepsilon_{n})}(\theta)\propto\pi(\theta)\int_{\mathcal{Y}^{n}}\mathds{1}\{\Delta(z_{1:n},y_{1:n})\leq\varepsilon_{n}\}\ \mu_{\theta}^{n}(dz_{1:n}),$
whose properties clearly depend on the chosen discrepancy $\Delta(\cdot,\cdot)$. Within summary-based abc, $\Delta(z_{1:n},y_{1:n})$ is a suitable distance, typically Euclidean, among summaries computed from the synthetic sample $z_{1:n}$ and the observed data $y_{1:n}$. However, recalling, e.g., Marin et al. (2014) and Frazier et al. (2018), the identification of summaries that do not lead to information loss is challenging for those complex models requiring abc implementations.
To overcome these challenges, abc literature has progressively moved towards the adoption of discrepancies $\mathcal{D}:\mathcal{P}(\mathcal{Y})\times\mathcal{P}(\mathcal{Y})\to[0,\infty]$ between the empirical distributions of the synthetic and observed data, that is
$\Delta(z_{1:n},y_{1:n})=\mathcal{D}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})=\mathcal{D}(n^{-1}\textstyle{\sum\nolimits_{i=1}^{n}}\delta_{z_{i}},n^{-1}\textstyle{\sum\nolimits_{i=1}^{n}}\delta_{y_{i}}),$
where $\delta_{x}$ is the Dirac delta at a generic $x\in\mathcal{Y}$. Popular examples are abc versions based on mmd, kl divergence, Wasserstein distance, energy statistic, Hellinger and Cramer–von Mises distances, and $\gamma$-divergence, whose limiting properties have been studied in Park, Jitkrittum and Sejdinovic (2016), Jiang, Wu and Wong (2018), Bernton et al. (2019), Nguyen et al. (2020), Frazier (2020) and Fujisawa et al. (2021) under different asymptotic regimes and with a common reliance on existence assumptions to ease the proofs. As a first step toward our unified and constructive theoretical framework, we shall emphasize that, although most of the above contributions treat the discrepancies separately, some of these choices share, in fact, a common origin. For example, mmd, Wasserstein distance and the energy statistic all belong to the ips class (Müller, 1997) in Definition 2.1. Such a class also includes summary-based distances.
Definition 2.1 (Integral probability semimetric — ips).
Let $\mathfrak{F}$ be a class of measurable functions $f:\mathcal{Y}\to\mathds{R}$. Then, the integral probability semimetric $\mathcal{D}_{\mathfrak{F}}$ between $\mu_{1},\mu_{2}\in\mathcal{P}(\mathcal{Y})$ is defined as
[…]
where $\mathfrak{F}$ is the pre-specified class of functions that characterizes $\mathcal{D}_{\mathfrak{F}}$.
Examples 2.22.4 show that routinely-employed discrepancies both in summary-free and summary-based abc (e.g., Park, Jitkrittum and Sejdinovic, 2016; Bernton et al., 2019; Nguyen et al., 2020; Drovandi and Frazier, 2022) are, in fact, ips with a known characterizing family $\mathfrak{F}$ that uniquely identifies each discrepancy.
Example 2.2.
(Wasserstein-1 distance). Let us consider the Lipschitz seminorm defined as $||f||_{L}:=\sup\{|f(x)-f(x^{\prime})|/\rho(x,x^{\prime}):x\neq x^{\prime}\mbox{ in }\mathcal{Y}\}$. If $\mathfrak{F}=\{f:||f||_{L}\leq 1\}$, then $\mathcal{D}_{\mathfrak{F}}$ coincides with the Kantorovich metric. Recalling the Kantorovich–Rubinstein theorem (e.g., Dudley, 2018), when $\mathcal{Y}$ is separable this metric coincides with the dual representation of the Wasserstein-1 distance, which is therefore recovered as an example of ips. Recall also that such a distance is defined on the space of probability measures having finite first moment (e.g., Sriperumbudur et al., 2012; Bernton et al., 2019).
Example 2.3.
(mmd and energy distance). Given a positive-definite kernel $k(\cdot,\cdot)$ on $\mathcal{Y}\times\mathcal{Y}$, let $\mathfrak{F}=\{f:||f||_{\mathcal{H}}\leq 1\}$, where $\mathcal{H}$ is the reproducing kernel Hilbert space corresponding to $k(\cdot,\cdot)$. Then $\mathcal{D}_{\mathfrak{F}}$ is the mmd (e.g., Muandet et al., 2017). When $k(\cdot,\cdot)$ is characteristic, then $\mathcal{D}_{\mathfrak{F}}$ is not only a semimetric, but also a metric — i.e., $\mathcal{D}_{\mathfrak{F}}(\mu_{1},\mu_{2})=0$ implies $\mu_{1}=\mu_{2}$. Relevant examples of routinely-implemented characteristic kernels in $\mathcal{Y}=\mathbb{R}^{d}$ are the Gaussian $\exp(-\|x-x^{\prime}\|^{2}/\sigma^{2})$ and Laplace $\exp(-\|x-x^{\prime}\|/\sigma)$ ones (Muandet et al., 2017). Note that mmd is also directly related to the energy distance, due to the correspondence among positive-definite kernels and negative-definite functions (Sejdinovic et al., 2013).
Example 2.4.
(Summary-based distances). Classical summary-based abc employs distances among a finite set of summaries $f_{1},\ldots,f_{K}$ (e.g., Drovandi and Frazier, 2022). Interestingly, these distances can be recovered as a special case of mmd with suitably-defined kernel and, hence, are guaranteed to belong to the ips class. In particular, for a pre-selected vector of summaries $f(x)=[f_{1}(x),\ldots,f_{K}(x)]\in\mathcal{H}=\mathbb{R}^{K}$ equipped with the standard Euclidean norm $\|f(x)\|_{\mathcal{H}}^{2}=\left<f(x),f(x)\right>=f^{2}_{1}(x)+\cdots+f^{2}_{K}(x)$, one can define the kernel $k(x,x^{\prime})=\left<f(x),f(x^{\prime})\right>$ to obtain classical summary-based abc recasted within the mmd framework. As a result, popular kernels, such as the Gaussian one, relying on an infinite dimensional feature space can be interpreted as a limiting version of summary-based abc.
Although Examples 2.22.4 characterize the most popular instances of ips discrepancies employed in abc, it shall be emphasized that other interesting semimetrics belong to the ips class (e.g., Sriperumbudur et al., 2012; Birrell et al., 2022). Two relevant ones are the total variation (tv) and Kolmogorov–Smirnov distances discussed in the Supplementary Material.
While abc based on discrepancies, such as those presented above, overcomes the need to pre-select summaries, it still requires the choice of a discrepancy. This motivates theory on the asymptotic properties of the induced abc posterior, under general discrepancies, for $n\rightarrow\infty$ and relevant thresholding schemes. Considering here a unified perspective which applies to the whole ips class, the current theory focuses both on fixed $\varepsilon_{n}=\varepsilon$ settings, and also on $\varepsilon_{n}\rightarrow\varepsilon^{*}$ regimes, where $\varepsilon^{*}=\inf\nolimits_{\theta\in\Theta}\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ is the lowest attainable discrepancy between the assumed model and the data generating process (Jiang, Wu and Wong, 2018; Bernton et al., 2019; Nguyen et al., 2020; Frazier, 2020; Frazier, Robert and Rousseau, 2020; Fujisawa et al., 2021). Under these settings, available theory investigates whether the $\varepsilon_{n}$-control on $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$ established by rejection-abc translates into meaningful convergence results, and upper bounds on the rates of concentration for the sequence of discrepancy-based abc posteriors around the true data generating process $\mu^{*}$, both under correctly specified models where $\mu^{*}=\mu_{\theta^{*}}$ for some $\theta^{*}\in\Theta$, and in misspecified contexts where $\mu^{*}$ does not belong to $\{\mu_{\theta}:\theta\in\Theta\subseteq\mathbb{R}^{p}\}$.
A common strategy to derive the aforementioned results is to study convergence of the empirical distributions $\hat{\mu}_{z_{1:n}}$ and $\hat{\mu}_{y_{1:n}}$ to the corresponding truths $\mu_{\theta}$ and $\mu^{*}$, in the chosen discrepancy $\mathcal{D}_{\mathfrak{F}}$, after noticing that, since $\mathcal{D}_{\mathfrak{F}}$ is a semimetric, we have
$\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*}).$ (1)
To this end, available theoretical studies pre-assume suitable convergence results for the empirical measures along with specific concentration inequalities regulated by non-explicit sequences of control functions; see Assumptions 1–2 in Proposition 3 of Bernton et al. (2019), and Assumptions A1–A2 in Nguyen et al. (2020). Alternatively, it is possible to directly require that $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\rightarrow\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ almost surely as $n\rightarrow\infty$ (see Jiang, Wu and Wong, 2018; Nguyen et al., 2020, Theorems 1 and 2, respectively). However, as highlighted by the same authors, these conditions (i) can be difficult to verify for several discrepancies, (ii) do not allow to assess whether some of these discrepancies can achieve convergence and concentration uniformly over $\mathcal{P}(\mathcal{Y})$, and (iii) often yield bounds which hinder an in-depth understanding of the factors regulating these limiting properties. In addition, when the above assumptions are not met, the asymptotic behavior of abc posteriors remains yet unexplored.
As a first step towards addressing the above issues, notice that the aforementioned results, when contextualized within the ips class, are inherently related to the richness of the known class of functions $\mathfrak{F}$ that identifies each $\mathcal{D}_{\mathfrak{F}}$; see Definition 2.1. Intuitively, if $\mathfrak{F}$ is too rich, it is always possible to find a function $f\in\mathfrak{F}$ yielding large discrepancies between the empiricals and the corresponding truths even when $\hat{\mu}_{z_{1:n}}$ and $\hat{\mu}_{y_{1:n}}$ are arbitrarily close to $\mu_{\theta}$ and $\mu^{*}$, respectively. Hence, $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})$ and $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$ will remain large with positive probability, making the triangle inequality in (1) of limited interest, since a low $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$ will not necessarily imply a small $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$. In fact, in this context, it is clearly even not guaranteed that $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$ will converge to $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$. This suggests that the limiting properties of the abc posterior are inherently related to the richness of $\mathfrak{F}$. In Section 3 we prove that this is the case when such a richness is measured through the notion of Rademacher complexity clarified in Definition 2.5; see also Chapter 4 in Wainwright (2019) for an introduction.
Definition 2.5 (Rademacher complexity).
Given a vector $x_{1:n}=(x_{1},\ldots,x_{n})$ of i.i.d. random variables with distribution $\mu\in\mathcal{P}(\mathcal{Y})$, the Rademacher complexity $\mathfrak{R}_{\mu,n}(\mathfrak{F})$ of a class $\mathfrak{F}$ of real-valued measurable functions is
$\mathfrak{R}_{\mu,n}(\mathfrak{F})=\mathbb{E}_{x_{1:n},\epsilon_{1:n}}\left[\sup\nolimits_{f\in\mathfrak{F}}\left|(1/n){\textstyle\sum\nolimits}_{i=1}^{n}\epsilon_{i}f(x_{i})\right|\right],$
where $\epsilon_{1},\dots,\epsilon_{n}$ denote i.i.d. Rademacher variables, that is $\mathbb{P}(\epsilon_{i}=1)=\mathbb{P}(\epsilon_{i}=-1)=1/2$.
As is clear from the above definition, high values of the Rademacher complexity $\mathfrak{R}_{\mu,n}(\mathfrak{F})$ mean that $\mathfrak{F}$ is rich enough to contain functions that can closely capture, on average, even the behavior of full-noise vectors. Hence, if $\mathfrak{R}_{\mu,n}(\mathfrak{F})$ is bounded away from zero for every $n$, then $\mathfrak{F}$ might be excessively rich for statistical purposes. Conversely, if $\mathfrak{R}_{\mu,n}(\mathfrak{F})$ goes to zero as $n$ diverges, then $\mathfrak{F}$ is a more parsimonious class. Lemma 2.6 formalizes this intuition.
Lemma 2.6 (Theorem 4.10 and Proposition 4.12 in Wainwright (2019)).
Let $x_{1:n}$ be i.i.d. from some distribution $\mu\in\mathcal{P}(\mathcal{Y})$. Then, for any $b$-uniformly bounded class $\mathfrak{F}$, i.e., any class $\mathfrak{F}$ of functions $f$ such that $\|f\|_{\infty}\leq b$, any integer $n\geq 1$ and scalar $\delta\geq 0$, it holds that
$\mathbb{P}_{x_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\leq 2\mathfrak{R}_{\mu,n}(\mathfrak{F})+\delta\right]\geq 1-\exp(-n\delta^{2}/2b^{2}),$ (2)
and $\mathbb{P}_{x_{1:n}}[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\geq\mathfrak{R}_{\mu,n}(\mathfrak{F})/2-\sup_{f\in\mathfrak{F}}|\mathbb{E}(f)|/2n^{1/2}-\delta]\geq 1-\exp(-n\delta^{2}/2b^{2})$.
Lemma 2.6 provides bounds for the probability that $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)$ takes values below a multiple and above a fraction of the Rademacher complexity of $\mathfrak{F}$. Recalling the previous discussion on concentration of abc posteriors, this result is crucial to study the convergence of both $\hat{\mu}_{z_{1:n}}$ and $\hat{\mu}_{y_{1:n}}$ to $\mu_{\theta}$ and $\mu^{*}$, respectively. Our theory in Section 3 proves that this is possible when $\mathfrak{R}_{n}(\mathfrak{F}):=\sup_{\mu\in\mathcal{P}(\mathcal{Y})}\mathfrak{R}_{\mu,n}(\mathfrak{F})$ goes to 00 as $n\rightarrow\infty$. According to Lemma 2.6, this is a necessary and sufficient condition for $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\rightarrow 0$ as $n\rightarrow\infty$, in $\mathbb{P}_{x_{1:n}}$–probability, uniformly over $\mathcal{P}(\mathcal{Y})$. See Appendix C in the Supplementary Material for extensions to non-i.i.d. settings.
3 Asymptotic properties of discrepancy-based abc posterior distributions
As anticipated in Section 2, the asymptotic results we derive leverage Lemma 2.6 to connect the properties of the abc posterior under an ips discrepancy $\mathcal{D}_{\mathfrak{F}}$ with the behavior of the Rademacher complexity for the associated family $\mathfrak{F}$. Section 3.1 clarifies the importance of this bridge by studying the limiting properties of the abc posterior when $\varepsilon_{n}=\varepsilon$ is fixed and $n\rightarrow\infty$, including in yet-unexplored situations where convergence guarantees for the empirical measures and for the discrepancy among such measures are not necessarily available. The theoretical results we derive in this case suggest that the use of discrepancies whose Rademacher complexity decays to zero in the limit is a sufficient condition to obtain strong convergence guarantees. On the basis of these findings, Section 3.2 focuses on the regime $\varepsilon_{n}\rightarrow\varepsilon^{*}=\inf\nolimits_{\theta\in\Theta}\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ as $n\rightarrow\infty$, and derives unified and informative concentration bounds that hold uniformly over $\mathcal{P}(\mathcal{Y})$ and are based on more constructive conditions than those employed in the available discrepancy-specific theory. More specifically, to prove our theoretical results in Sections 3.13.2, we rely on some or all of the following assumptions:
  1. (I)
    the observed data $y_{1:n}$ are i.i.d. from a data generating process $\mu^{*}$;
  2. (II)
    there exist $L,c_{\pi}\in\mathds{R}^{+}$ such that, for $\bar{\varepsilon}$ small, $\pi(\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon^{*}+\bar{\varepsilon}\})\geq c_{\pi}\bar{\varepsilon}^{L}$;
  3. (III)
    $\mathfrak{F}$ is $b$-uniformly bounded; i.e., there exists $b\in\mathds{R}^{+}$ such that $||f||_{\infty}{\leq b}$ for any $f\in\mathfrak{F}$;
  4. (IV)
    $\mathfrak{R}_{n}(\mathfrak{F}):=\sup_{\mu\in\mathcal{P}(\mathcal{Y})}\mathfrak{R}_{\mu,n}(\mathfrak{F})$ goes to zero as $n\to\infty$.
Condition (I) is the only assumption made on the data generating process and is present in, e.g., Nguyen et al. (2020) and in the supplementary materials of Bernton et al. (2019). Although the theory we derive in Appendix C of the Supplementary Material relaxes (I) to study convergence and concentration also beyond the i.i.d. context, it shall be emphasized that some of the assumptions considered in the literature may not hold even in i.i.d. settings. Hence, an improved understanding of the abc properties under (I) is crucial to clarify the range of applicability and potential limitations of available existence theory under more complex, possibly non-i.i.d. regimes. In fact, as shown within Section 3.1, certain discrepancies may yield posteriors that are either not well-defined in the limit or lack strong convergence guarantees even in i.i.d. settings. Assumption (II) is standard in Bayesian asymptotics and abc theory (Bernton et al., 2019; Nguyen et al., 2020; Frazier, 2020), and requires that sufficient prior mass is placed on those parameter values yielding models $\mu_{\theta}$ close to $\mu^{*}$, under $\mathcal{D}_{\mathfrak{F}}$. Finally, (III)–(IV) provide conditions on $\mathfrak{F}$ to ensure that $\mathcal{D}_{\mathfrak{F}}$ yields meaningful concentration bounds, leveraging Lemma 2.6. Crucially, (III)–(IV) are made directly on the user-selected discrepancy $\mathcal{D}_{\mathfrak{F}}$, rather than on the data generating process or on the statistical model, thereby providing general constructive conditions that are more realistic to verify in practice compared to regularity assumptions on the unknown $\mu^{*}$ or on $\{\mu_{\theta}:\theta\in\Theta\subseteq\mathbb{R}^{p}\}$, while guiding the choice of $\mathcal{D}_{\mathfrak{F}}$.
Notice that (III)–(IV) depend, in principle, on $\mathcal{P}(\mathcal{Y})$. Nonetheless, as clarified within Section 3.3, available bounds on the Rademacher complexity ensure that these two conditions hold for the relevant ips discrepancies in Examples 2.22.4 under either no constraints on $\mathcal{P}(\mathcal{Y})$ — thereby circumventing such a dependence — or for suitable spaces $\mathcal{Y}$ (e.g., bounded ones). Unlike for constraints on the unknown $\mu^{*}$ and on $\{\mu_{\theta}:\theta\in\Theta\subseteq\mathbb{R}^{p}\}$, conditions on $\mathcal{Y}$ can be verified in practice; see Examples 3.53.7, and refer to the Supplementary Material for the analysis of other relevant ips discrepancies. Moreover, Assumptions (III) and (IV) are generally not stronger than those implicit in the current theory on discrepancy-based abc, and are inherently related with the notion of uniform Glivenko–Cantelli classes (see e.g., Wainwright, 2019, Chapter 4), arguably a minimal sufficient requirement for establishing convergence and concentration properties that, unlike those currently derived, hold uniformly over $\mathcal{P}(\mathcal{Y})$. This latter point is further clarified in Propositions 4.3 and 4.4, which state concentration results for mmd with unbounded kernel in $\mathbb{R}^{d}$ (which also includes abc with unbounded summary statistics) and Wasserstein-1 distance in $\mathbb{R}^{d}$, respectively. Although these two discrepancies meet (III)–(IV) for suitable constraints on $\mathcal{Y}$, in general unbounded $\mathcal{Y}$ settings these conditions are no more guaranteed for such discrepancies. In this case, Propositions 4.3 and 4.4 clarify that concentration properties can still be derived, but at the expense of additional constraints on $\mu^{*}$ and $\{\mu_{\theta}:\theta\in\Theta\subseteq\mathbb{R}^{p}\}$, which yield results that hold uniformly only in a suitable subset of $\mathcal{P}(\mathcal{Y})$, rather than over the entire $\mathcal{P}(\mathcal{Y})$. It shall be also emphasized that the convergence theory derived in Section 3.1 does not necessarily require a vanishing Rademacher complexity to obtain informative results. However, it also clarifies that, when this condition is verified, the limiting pseudo-posterior inherits the same threshold-control on $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ as the one established by rejection abc on $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$.
As anticipated above, Sections 3.13.2 consider the two scenarios where the sample size $n$ diverges to infinity, while the abc threshold $\varepsilon_{n}$ is either fixed at $\varepsilon$ or progressively shrinks to $\varepsilon^{*}=\inf\nolimits_{\theta\in\Theta}\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$. If the model is well-specified, i.e., $\mu^{*}=\mu_{\theta^{*}}$ for some, possibly not unique, $\theta^{*}\in\Theta$, then $\varepsilon^{*}=0$. Although the regime characterized by $\varepsilon_{n}\to 0$ with fixed $n$ is also of interest, this setting can be readily addressed under the unified ips class via a direct adaptation of available results in, e.g., Jiang, Wu and Wong (2018) and Bernton et al. (2019); see also Miller and Dunson (2019) for results in the context of coarsened posteriors under the regime with $n\to\infty$ and fixed $\varepsilon_{n}=\varepsilon$, related to those in Jiang, Wu and Wong (2018).
3.1 Limiting behavior with fixed $\varepsilon_{n}=\varepsilon$ and $n\rightarrow\infty$
Current abc theory for the setting $\varepsilon_{n}=\varepsilon$ fixed and $n\rightarrow\infty$, studies convergence of the discrepancy-based abc posterior
${\pi^{(\varepsilon)}_{n}(\theta)}\propto\pi(\theta)\int_{\mathcal{Y}^{n}}\mathds{1}\{\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon\}\mu_{\theta}^{n}(dz_{1:n}),$
under the key assumption that $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\rightarrow\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ almost surely as $n\rightarrow\infty$ (see, e.g., Jiang, Wu and Wong, 2018, Theorem 1). When such a condition is met, it can be shown that $\lim_{n\rightarrow\infty}{\pi^{(\varepsilon)}_{n}(\theta)}\propto\pi(\theta)\mathds{1}\{\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon\}$. This means that the limiting abc posterior coincides with the prior constrained within the support $\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon\}$, where $\varepsilon$ is the same threshold employed for $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$.
Although this is a key result, verifying the convergence assumption behind the above finding can be difficult for several discrepancies (e.g., Jiang, Wu and Wong, 2018). More crucially, such theory fails to provide informative results in those situations where $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$ is not guaranteed to converge to $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$. In fact, under this setting, the asymptotic properties of discrepancy-based abc posteriors have been overlooked, and it is not even clear whether a well-defined ${\pi^{(\varepsilon)}_{n}}$ exists in the limit. Indeed, for specific choices of the discrepancy and threshold, the abc posterior may not be well-defined even for fixed $n$. For example, when $\mathcal{D}_{\mathfrak{F}}$ is the tv distance (see Appendix A in the Supplementary Material), then $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})=1$, almost surely, whenever $z_{1:n}$ and $y_{1:n}$ are from two continuous distributions $\mu_{\theta}$ and $\mu^{*}$, respectively. Hence, for any $\varepsilon<1$, the abc posterior is not defined, even if $\mu_{\theta}$ and $\mu^{*}$ are not mutually singular. Note that, as clarified in Appendix A, the Rademacher complexity of the tv distance is always bounded away from zero in such a setting. In this case and, more generally, for ips discrepancies whose $\mathfrak{R}_{n}(\mathfrak{F})$ does not vanish to zero, Lemma 2.6 together with the triangle inequality in (1) still allow to derive new informative upper and lower bounds for limiting acceptance probabilities which further guarantee that $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ can be bounded from above, for a sufficiently large $n$, by $\varepsilon+4\mathfrak{R}_{n}(\mathfrak{F})$, even without assuming that $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\rightarrow\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ almost surely for $n\rightarrow\infty$, as in the current convergence theory. Theorem 3.1 formalizes this intuition for the whole class of abc posteriors arising from discrepancies in the ips family.
Theorem 3.1.
Let $\mathcal{D}_{\mathfrak{F}}$ be an ips as in Definition 2.1. Moreover, assume (I), (III), and let $c_{\mathfrak{F}}=4\lim\sup\mathfrak{R}_{n}(\mathfrak{F})$. Then, the acceptance probability $p_{n}=\int_{\Theta}\int_{\mathcal{Y}^{n}}\mathds{1}\{\mathcal{D}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon\}\mu_{\theta}^{n}(dz_{1:n})\pi(d\theta)$ of rejection abc with discrepancy $\mathcal{D}_{\mathfrak{F}}$ satisfies, for any fixed $\varepsilon>0$,
$\qquad\pi\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon-c_{\mathfrak{F}}\}-e_{n}\leq p_{n}\leq\pi\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon+c_{\mathfrak{F}}\}+e_{n},$ (3)
almost surely with respect to $y_{1:n}\stackrel{{\scriptstyle\mbox{\scriptsize i.i.d.}}}{{\sim}}\mu^{*}$, as $n\rightarrow\infty$, where $e_{n}=\exp(-\sqrt{n}/(2b^{2}))=o(1)$.
In particular, whenever $\varepsilon>\tilde{\varepsilon}+c_{\mathfrak{F}}$, with $\tilde{\varepsilon}=\inf\{\epsilon>0:{\pi\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\epsilon\}}>0\}$, the probability $p_{n}$ is bounded away from 00, for $n$ large enough. This implies that, for every fixed $\varepsilon>\tilde{\varepsilon}+c_{\mathfrak{F}}$, the abc posterior is always well-defined for a large-enough $n$ and its support is almost surely included asymptotically in the set $\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon+4\mathfrak{R}_{n}(\mathfrak{F})\}$. Namely,
$\pi_{n}^{(\varepsilon)}\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon+4\mathfrak{R}_{n}(\mathfrak{F})\}\rightarrow 1,$
for any fixed $\varepsilon>\tilde{\varepsilon}+c_{\mathfrak{F}}$, almost surely with respect to $y_{1:n}\stackrel{{\scriptstyle\mbox{\scriptsize i.i.d.}}}{{\sim}}\mu^{*}$, as $n\rightarrow\infty$.
Theorem 3.1 clarifies that, even when relaxing the assumptions behind current theory, the Rademacher complexity framework still allows to obtain guarantees on the acceptance probabilities, existence and limiting support of the abc posterior. Such a relaxation highlights the key role of the richness of $\mathfrak{F}$ in driving these properties. As previously discussed, the higher is $\mathfrak{R}_{n}(\mathfrak{F})$ the lower are the guarantees that $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$ is close enough to $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ in the large data limit. In fact, recalling Section 2, when $\mathfrak{R}_{n}(\mathfrak{F})$ does not vanish with $n$ there are no guarantees that the convergence of $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$ to $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ assumed in the literature is verified. Nonetheless, the inequalities in Lemma 2.6 still yield informative results that translate these arguments into an inflation of the superset containing the support of the limiting abc posterior, and in a more conservative choice for the threshold $\varepsilon$ to ensure that such a posterior is well-defined for large $n$. Notice that in Theorem 3.1 the dependence on the dimension of the parameter and data space is embedded within $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$; see Equation (6) in the Appendix of Chérief-Abdellatif and Alquier (2022) for an example making this dependence explicit.
Recall that, for the abc posterior to be well-defined in the limit (with a fixed threshold $\varepsilon$), the acceptance probability $p_{n}$ needs to be bounded away from zero. According to Theorem 3.1, as $n\rightarrow\infty$ this is guaranteed whenever $\varepsilon>\tilde{\varepsilon}+c_{\mathfrak{F}}$. For instance, recalling Section 2, in misspecified models $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\geq\varepsilon^{*}$. As a consequence, $\varepsilon>c_{\mathfrak{F}}$ is not sufficient to guarantee that $\pi\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon-c_{\mathfrak{F}}\}$ in (3) is bounded away from zero (in fact, when $c_{\mathfrak{F}}<\varepsilon<c_{\mathfrak{F}}+\varepsilon^{*}$ the lower bound in (3) approaches zero from below). To achieve this result, $\varepsilon-c_{\mathfrak{F}}$ needs to define a ball with radius not lower than the one to which the prior is already guaranteed to assign strictly positive mass. This implies $\varepsilon>\tilde{\varepsilon}+c_{\mathfrak{F}}$, with $c_{\mathfrak{F}}=4\lim\sup\mathfrak{R}_{n}(\mathfrak{F})$.
The results in Theorem 3.1 further clarify that, when $\mathfrak{R}_{n}(\mathfrak{F})\rightarrow 0$ as $n\rightarrow\infty$, stronger and uniform convergence statements can be derived, both in correctly-specified and in misspecified models. More specifically, in this setting, the acceptance probability $p_{n}$ in Theorem 3.1 approaches $\pi\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon\}$ in the limit. Moreover, the abc posterior is well-defined for any $\varepsilon>\tilde{\varepsilon}$ and its asymptotic support has the same $\varepsilon$-control as the one established by the abc procedure on the discrepancy among the empirical distributions. These results allow to establish the almost sure convergence of the abc posterior stated in Corollary 3.2.
Corollary 3.2.
Under the conditions in Theorem 3.1, if also $\mathfrak{R}_{n}(\mathfrak{F})\rightarrow 0$ as $n\rightarrow\infty$ (i.e., Assumption (IV) is satisfied), then, for any fixed $\varepsilon>\tilde{\varepsilon}$, it holds that
$\pi_{n}^{(\varepsilon)}(\theta)\rightarrow\pi(\theta\mid\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon)\propto\pi(\theta)\mathds{1}\left\{\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon\right\},$ (4)
almost surely with respect to $y_{1:n}\stackrel{{\scriptstyle\mbox{\scriptsize i.i.d.}}}{{\sim}}\mu^{*}$, as $n\rightarrow\infty$.
Corollary 3.2 clarifies that to obtain within the whole ips class a convergence result in line with those stated in the current theory it is not necessary to assume that $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$ converges almost surely to $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$. In fact, it is sufficient to verify that $\mathfrak{R}_{n}(\mathfrak{F})\rightarrow 0$ as $n\rightarrow\infty$. Crucially, this condition is imposed directly on the selected discrepancy within the ips class, rather than on the model or the underlying data generating process and, therefore, can be constructively verified for all the discrepancies in Examples 2.22.4. For instance, while Jiang, Wu and Wong (2018) is unable to provide conclusive guidelines on whether (4) holds under Wasserstein distance and mmd, our Corollary 3.2 can be directly applied to prove convergence for both these divergences, leveraging the results and discussions in Section 3.3.
Corollary 3.2 accounts also for misspecified models. In fact, the limiting pseudo-posterior in (4) is well-defined only when $\varepsilon>\tilde{\varepsilon}$. For example, when $\mu^{*}$ is not within $\{\mu_{\theta}:\theta\in\Theta\subseteq\mathbb{R}^{p}\}$, then $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\geq\varepsilon^{*}$. Hence, for any $\varepsilon\leq\varepsilon^{*}$, we have that $\mathds{1}\{\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon\}=0$ for any $\theta\in\Theta$ and the limiting pseudo-posterior is not well-defined.
Motivated by the convergence result in Corollary 3.2, the theory in Section 3.2 provides an in-depth study of the concentration properties for the abc posterior when the tolerance threshold progressively shrinks. As is clear from (4), employing a vanishing threshold might guarantee that, in the limit, the abc posterior concentrates all its mass around $\mu^{*}$, in correctly specified settings, or around the distribution $\mu_{\theta^{*}}$ closest to $\mu^{*}$, among those in $\{\mu_{\theta}:\theta\in\Theta\subseteq\mathbb{R}^{p}\}$ for misspecified regimes. These results and the associated rates of concentration are provided in Section 3.2 by leveraging, again, Rademacher complexity theory.
3.2 Concentration as $\varepsilon_{n}\rightarrow\varepsilon^{*}$ and $n\rightarrow\infty$
Theorem 3.3 states our main concentration result. As in related abc theory (see e.g., Proposition 3 in Bernton et al. (2019) and Nguyen et al. (2020)), we also leverage the triangle inequality (1). However, we crucially avoid pre-assuming convergence of $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$, and we do not rely on concentration inequalities for $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}}{,}\mu_{\theta})$ regulated by non-explicit sequences of control functions $c_{n}(\theta)f_{n}(\bar{\varepsilon}_{n})$ (Bernton et al., 2019) and $c_{n}(\theta)s_{n}(\bar{\varepsilon}_{n})$ (Nguyen et al., 2020). Rather, we leverage Lemma 2.6 under the proposed Rademacher complexity framework to obtain more direct and informative results, that are also guaranteed to hold uniformly over $\mathcal{P}(\mathcal{Y})$. Indeed, Lemma 2.6 ensures that both $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$ and $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})$ exceed $2\mathfrak{R}_{n}(\mathfrak{F})$ with vanishing probability. Therefore, when $\mathfrak{R}_{n}(\mathfrak{F})\rightarrow 0$, we have that $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\approx\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$, by combining (1) with
$\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\geq-\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})-\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*}).$
This means that, if $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$ is small, then $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ is also small with a high probability. This clarifies the importance of Assumption (IV), which is further supported by the fact that, if $\mathfrak{R}_{n}(\mathfrak{F})$ does not shrink to zero, by the second inequality in Lemma 2.6, the two empirical measures $\hat{\mu}_{y_{1:n}}$ and $\hat{\mu}_{z_{1:n}}$ do not converge to $\mu^{*}$ and $\mu_{\theta}$, respectively, and hence there is no guarantee that a small $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$ would imply a vanishing $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$.
Theorem 3.3.
Let $\bar{\varepsilon}_{n}\to 0$ when $n\to\infty$, with $n\bar{\varepsilon}^{2}_{n}\to\infty$ and $\bar{\varepsilon}_{n}/\mathfrak{R}_{n}(\mathfrak{F})\to\infty$. Then, if $\mathcal{D}_{\mathfrak{F}}$ is from the ips class in Definition 2.1 and Assumptions (I)–(IV) hold, the abc posterior with threshold $\varepsilon_{n}=\varepsilon^{*}+\bar{\varepsilon}_{n}$ satisfies
${\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}\Bigl{(}\Bigl{\{}\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})}>{\varepsilon^{*}+\frac{4\bar{\varepsilon}_{n}}{3}+2\mathfrak{R}_{n}(\mathfrak{F})+\Bigl{(}\frac{2b^{2}}{n}\log\frac{n}{\bar{\varepsilon}_{n}^{L}}\Bigr{)}^{1/2}}\Bigr{\}}\Bigr{)}\leq\frac{2\cdot 3^{L}}{c_{\pi}n},$
with $\mathbb{P}_{y_{1:n}}$–probability going to $1$ as $n\rightarrow\infty.$
The proof of Theorem 3.3 can be found in Appendix D of the Supplementary Material and follows similar arguments considered to establish the concentration results in Bernton et al. (2019) and Nguyen et al. (2020), which, in turn, extend those in Frazier et al. (2018). However, as mentioned above, those proofs are specific to a single discrepancy, pre-assume the convergence of $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$, and rely on concentration inequalities for $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})$ that depend on non-explicit sequences of control functions. Theorem 3.3 overcomes these issues and proves a unified theory based on the single concentration inequality in Lemma 2.6. This yields technical differences in the proof and, more importantly, it introduces a novel and broadly-impactful perspective for the analysis of concentration properties of discrepancy-based abc posteriors.
Note that in Theorem 3.3 the constant $b$ can be typically set equal to $1$ either by definition or upon normalization of the class of $b$-uniformly bounded functions. Moreover, as clarified in the Supplementary Material, Theorem 3.3 also holds when replacing $n/\bar{\varepsilon}_{n}^{L}$ and $c_{\pi}n$ with $M_{n}/\bar{\varepsilon}_{n}^{L}$ and $c_{\pi}M_{n}$, respectively, for any sequence $M_{n}>1$. Nonetheless, to ensure concentration, it suffices to let $M_{n}=n$. In such a case, the quantities $34\bar{\varepsilon}_{n}/$ 3, $2\mathfrak{R}_{n}(\mathfrak{F})$, $[(2b^{2}/n)\log(n/\bar{\varepsilon}_{n}^{L})]^{1/2}$ and $2\cdot 3^{L}/c_{\pi}n$ converge to 00 as $n\rightarrow\infty$, under the settings of Theorem 3.3. This implies that the abc posterior asymptotically concentrates around those $\theta$ values yielding a $\mu_{\theta}$ within discrepancy $\varepsilon^{*}$ from $\mu^{*}$. Appendix B in the Supplementary Material translates these concentration results from the space of distributions to the space of parameters.
In contrast to available theory, the concentration result in Theorem 3.3 holds uniformly over $\mathcal{P}(\mathcal{Y})$, and replaces the currently-employed non-explicit control functions with a known and well-studied quantity, i.e., the Rademacher complexity. Notice that, although Theorem 3.3 provides a unified statement for all the ips discrepancies, the bound we derive depends on $\mathfrak{R}_{n}(\mathfrak{F})$, which is specific to each discrepancy $\mathcal{D}_{\mathfrak{F}}$ and plays a fundamental role in controlling the rate of concentration of the abc posterior. In particular, to make the bound as tight as possible, we must choose an $\bar{\varepsilon}_{n}$ such that $34\bar{\varepsilon}_{n}/$ 3 and $[(2b^{2}/n)\log(n/\bar{\varepsilon}_{n}^{L})]^{1/2}$ are of the same order. By neglecting all the terms in $\log\log n$, such a choice leads to setting $\bar{\varepsilon}_{n}$ of the order $[\log(n)/n]^{1/2}$. In this case, the constraint $\bar{\varepsilon}_{n}/\mathfrak{R}_{n}(\mathfrak{F})\to\infty$ may not be satisfied, thereby requiring a larger $\bar{\varepsilon}_{n}$, such as $\mathfrak{R}_{n}(\mathfrak{F})\log\log(n)$. Summarizing, when $\bar{\varepsilon}_{n}=\max\{[\log(n)/n]^{1/2},\mathfrak{R}_{n}(\mathfrak{F})\log\log(n)\}$ it follows, under the conditions of Theorem 3.3, that
${\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}(\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})}>\varepsilon^{*}+\mathcal{O}(\bar{\varepsilon}_{n})\})\leq 2\cdot 3^{L}/(c_{\pi}n).$
Notice that the conditions $n\bar{\varepsilon}^{2}_{n}\to\infty$ and $\bar{\varepsilon}_{n}/\mathfrak{R}_{n}(\mathfrak{F})\to\infty$ do not allow to set $\bar{\varepsilon}_{n}<1/\sqrt{n}$. Although this regime is of interest, we are not aware of explicit results in the discrepancy-based abc literature for such a setting. In fact, available studies rely on similar restrictions for some sequence of functions $f_{n}(\bar{\varepsilon}_{n})$ which is not made explicit except for specific examples that still point toward setting $\bar{\varepsilon}_{n}=[\log(n)/n]^{1/2}>1/\sqrt{n}$ (e.g., Bernton et al., 2019, supplementary materials). Faster rates for the abc threshold have been considered by Li and Fearnhead (2018) in the context of summary-based abc, but with a substantially different theoretical focus and asymptotic regime relative to the one considered here. Recalling, e.g., Theorem 1 in Frazier et al. (2018), within our theory setting the rate $1/\sqrt{n}$ (up to log terms) cannot be improved for summary-based abc; see also Section 4.1 for additional discussion. It shall be also emphasized that $\bar{\varepsilon}_{n}<1/\sqrt{n}$ would not yield a faster concentration rate in Theorem 3.3 because of the terms $2\mathfrak{R}_{n}(\mathfrak{F})$ and $[(2b^{2}/n)\log(n/\bar{\varepsilon}_{n}^{L})]^{1/2}$ in the bound.
Theorem 3.3 holds under both well-specified and misspecified models. In the former case, $\mu^{*}=\mu_{\theta^{*}}$ for some $\theta^{*}\in\Theta$. Therefore, $\varepsilon^{*}=0$ and the abc posterior concentrates around those $\theta$ yielding $\mu_{\theta}=\mu^{*}$. Conversely, when the model is misspecified, the abc posterior concentrates on those $\theta$ yielding a $\mu_{\theta}$ within discrepancy $\varepsilon^{*}$ from $\mu^{*}$. Since $\varepsilon^{*}=\inf\nolimits_{\theta\in\Theta}\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$, the abc posterior will concentrate on those $\theta$ such that $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ is arbitrarily close to $\varepsilon^{*}$. Example 3.4 provides an explicit bound for $\varepsilon^{*}$ in a simple yet noteworthy class of misspecified models.
Example 3.4.
(Huber contamination model). In this model, with probability $1-\alpha_{n}$, the data are from a distribution $\mu_{\theta^{*}}$ belonging to the assumed model $\{\mu_{\theta}:\theta\in\Theta\subseteq\mathbb{R}^{p}\}$, while with probability $\alpha_{n}$ arise from a contaminating distribution $\mu_{C}$. Therefore, the data generating process is $\mu^{*}=(1-\alpha_{n})\mu_{\theta^{*}}+\alpha_{n}\mu_{C},$ with $\alpha_{n}\in[0,1)$ controlling the amount of contamination. In such a context, Definition 2.1 and Assumption (III) imply
$\varepsilon^{*}\leq\mathcal{D}_{\mathfrak{F}}(\mu_{\theta^{*}},\mu^{*})=\mathcal{D}_{\mathfrak{F}}(\mu_{\theta^{*}},(1-\alpha_{n})\mu_{\theta^{*}}+\alpha_{n}\mu_{C})=\alpha_{n}\mathcal{D}_{\mathfrak{F}}(\mu_{\theta^{*}},\mu_{C})\leq 2b\alpha_{n}.$
By plugging this bound into Theorem 3.3, we can obtain the same statement with $\varepsilon^{*}$ replaced by $2b\alpha_{n}$, where $\alpha_{n}\in[0,1)$ is the amount of contamination. This means that the abc posterior asymptotically contracts in a neighborhood of the contaminated model $\mu^{*}$ of radius at most $2b\alpha_{n}$. The previous bound on $\varepsilon^{*}$ combined with a triangle inequality also implies
$\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\geq\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu_{\theta^{*}})-\mathcal{D}_{\mathfrak{F}}(\mu_{\theta^{*}},\mu^{*})\geq\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu_{\theta^{*}})-2b\alpha_{n}.$
Therefore, replacing $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ with $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu_{\theta^{*}})$ guarantees concentration also around the uncontaminated model $\mu_{\theta^{*}}$, with $\varepsilon^{*}$ replaced by $2b\alpha_{n}+2b\alpha_{n}=4b\alpha_{n}$, thus ensuring robustness to Huber contamination.
3.3 Validity of the assumptions
The main theorems in Section 3.13.2 (i.e., Theorems 3.1 and 3.3) leverage Assumptions (I)–(IV). As anticipated in Section 3, Assumption (I) is useful to formally clarify the range of applicability along with the possible limitations of the current existence theory even in simple i.i.d. settings, and it will be relaxed in Appendix C of the Supplementary Material to study uniform convergence and concentration of discrepancy-based abc posteriors beyond the i.i.d. context. Assumption (II) is not specific to our framework. Rather, it defines a standard minimal requirement routinely employed in Bayesian asymptotics and abc theory (e.g., Bernton et al., 2019; Nguyen et al., 2020; Frazier, 2020). Conversely, Assumptions (III)–(IV) provide sufficient conditions on the ips discrepancy $\mathcal{D}_{\mathfrak{F}}$ to obtain guarantees of uniform convergence and concentration under the proposed Rademacher complexity perspective. These two conditions essentially replace the pre-assumed convergence for $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$ and the non-explicit bounds within the concentration inequalities for $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})$ leveraged by the current discrepancy-specific theory. While these latter assumptions may implicitly require regularity conditions on $\{\mu_{\theta}:\theta\in\Theta\subseteq\mathbb{R}^{p}\}$ and on the unknown data generating process $\mu^{*}$, Assumptions (III)–(IV) are made directly on the user-selected discrepancy $\mathcal{D}_{\mathfrak{F}}$. Leveraging the available bounds on the Rademacher complexity, these two conditions crucially allow to state results that hold uniformly over $\mathcal{P}(\mathcal{Y})$.
Notice that, to derive these uniform convergence and concentration properties, an arguably minimal sufficient requirement is that $\mathfrak{F}$ is a uniform Glivenko–Cantelli class (see e.g., Dudley, Giné and Zinn, 1991, Proposition 10), i.e., for $x_{1:n}$ i.i.d. from $\mu\in\mathcal{P}(\mathcal{Y})$,
[…]
The above connection clarifies that Assumptions (III)–(IV) are arguably at the core of the uniform convergence and concentration properties of discrepancy-based abc posteriors. Moreover, although (5) is inherently related to (III)–(IV), such a uniform Glivenko–Cantelli property only states a convergence in probability result which can be crucially refined through the notion of Rademacher complexity under the more precise concentration inequalities in Lemma 2.6. Recalling the theoretical results in Sections 3.1 and 3.2, this allows not only to state convergence and concentration of specific abc posteriors, but also to clarify the factors governing these limiting properties and possibly derive the associated rates.
As clarified in Examples 3.53.7, Assumptions (III)–(IV) can be generally verified for the key ips discrepancies presented in Examples 2.22.4 leveraging known upper bounds on the Rademacher complexity, along with connections between this measure and other well-studied quantities in statistical learning theory. See, in particular, Chapter 4.3 of Wainwright (2019) for an overview of several useful techniques for upper-bounding the Rademacher complexity via, e.g., the notion of polynomial discrimination and vc dimension. The validity of (III)–(IV) for two other instances of the ips class (i.e., the total variation distance and the Kolmogorov–Smirnov distance) is discussed in Appendix A within the Supplementary Material. Notice that, albeit interesting, these two discrepancies are less common in abc implementations relative to Wasserstein distance, mmd and summary-based distances.
Example 3.5.
(Wasserstein-1 distance). When $\mathcal{Y}$ is a bounded subset of $\mathbb{R}^{d}$, Assumptions (III)–(IV) hold without further constraints under the Wasserstein-1 distance. In particular, (III) follows immediately from the definition $\mathfrak{F}=\{f:||f||_{L}\leq 1\}$, together with the fact that the diameter of $\mathcal{Y}$ is finite in this case (e.g., Villani, 2021, Remark 1.15). Assumption (IV) is instead a direct consequence of the bounds in Sriperumbudur et al. (2012). Although it would be desirable to remove such a constraint on $\mathcal{Y}$, it shall be emphasized that this condition is ubiquitous in state-of-the-art concentration results of empirical measures, under the Wasserstein distance, that are guaranteed to hold uniformly over $\mathcal{P}(\mathcal{Y})$ (e.g., Talagrand, 1994; Sriperumbudur et al., 2012; Ramdas, Trillos and Cuturi, 2017; Weed and Bach, 2019). One possibility to preserve (III)–(IV) under the Wasserstein-1 distance beyond bounded $\mathcal{Y}$ is to consider a variable transformation via a monotone function $g(\cdot)$ (e.g., logistic transform) mapping from $\mathcal{Y}=\mathbb{R}^{d}$ to a bounded subset of $\mathbb{R}^{d}$. In the original unbounded $\mathcal{Y}$, this transformation induces a Wasserstein-1 distance based on a bounded $\bar{\rho}(x,x^{\prime})=\rho(g(x),g(x^{\prime}))$. As such, when $\mathcal{Y}$ is not a bounded subset of $\mathbb{R}^{d}$, defining $\mathfrak{F}$ as the class of Lipschitz functions with respect to $\bar{\rho}(x,x^{\prime})=\rho(g(x),g(x^{\prime}))$ satisfies (III) and (IV). However, this requires care and further research on how $g(\cdot)$ affects the properties of the discrepancy. Alternatively, as shown in Proposition 4.4 within Section 4.2, concentration results for Wasserstein-1 distance in unbounded $\mathcal{Y}\subseteq\mathbb{R}^{d}$ can be still derived, but at the expense of some restrictions on $\mu\in\mathcal{P}(\mathcal{Y})$.
Example 3.6.
(mmd). The properties of mmd inherently depend on the selected kernel $k(\cdot,\cdot)$. This is evident from the inequalities $\mathfrak{R}_{\mu,n}(\mathfrak{F})\leq[\mathbb{E}_{x}k(x,x)/n]^{1/2}$, with $x\sim\mu$ (see e.g., Lemma 22 in Bartlett and Mendelson, 2002), and $|f(x)|\leq[k(x,x)]^{1/2}||f||_{\mathcal{H}}$ for every $x\in\mathcal{Y}$ (see e.g., equation 16 in Hofmann, Schölkopf and Smola, 2008). By these two inequalities, all bounded kernels, including, e.g., the commonly-implemented Gaussian $\exp(-\|x-x^{\prime}\|^{2}/\sigma^{2})$ and the Laplace $\exp(-\|x-x^{\prime}\|/\sigma)$ ones, ensure that Assumptions (III) and (IV) are always met without requiring additional regularity conditions on $\mu\in\mathcal{P}(\mathcal{Y})$, nor constraints on $\mathcal{Y}$. Instead, when $k(\cdot,\cdot)$ is unbounded, the inequality $\mathfrak{R}_{\mu,n}(\mathfrak{F})\leq[\mathbb{E}_{x}k(x,x)/n]^{1/2}$ is only informative for those $\mu$ such that $\mathbb{E}_{x}k(x,x)<\infty$, with $x\sim\mu$, whereas the bound on $|f(x)|$ does not guarantee that $\mathfrak{F}$ is a uniformly bounded class in general, unless additional conditions are made. Due to the relevance of these mmd instances and the direct connections with the classical summary-based abc implementations leveraging unbounded summaries, Proposition 4.3 in Section 4.1 derives specific theory proving that concentration results similar to those in Theorem 3.3 can be derived, under different assumptions, including conditions on $\mu\in\mathcal{P}(\mathcal{Y})$, also for unbounded kernels.
Example 3.7.
(Summary-based distance). As discussed in Example 2.4, classical abc implementations relying on a finite set of summaries $f_{1},\ldots,f_{K}$ with $K<\infty$, can be seen as a special case of mmd by letting $f(x)=[f_{1}(x),\ldots,f_{K}(x)]$ and $k(x,x^{\prime})=\left<f(x),f(x^{\prime})\right>$. Leveraging this bridge and the results for mmd discussed in Example 3.6, it is clear that, if $\sup_{x\in\mathcal{Y}}\left<f(x),f(x)\right>$ is finite — i.e., the induced kernel is bounded — then (III) and (IV) are satisfied without requiring regularity conditions for $\mu$ or constrains on $\mathcal{Y}$. While this result clarifies that abc with bounded summaries achieves uniform convergence and concentration, classical abc implementations often employ unbounded summaries, such as moments, i.e., $f(x)=[x,x^{2}\ldots,x^{K}]$. In this case (III) is not satisfied. In fact, recalling again Example 2.4, such a setting is a special case of mmd with unbounded kernel and, hence, lacks guarantees that (III)–(IV) hold, unless further conditions are imposed, e.g., on $\mathcal{Y}$. Nonetheless, this connection also clarifies that the concentration theory we derive in Proposition 4.3 for mmd with unbounded kernels directly applies to classical abc with unbounded summaries.
Examples 3.53.7 show that (III)–(IV) can be realistically verified for the key instances of the ips class presented in Section 2, and generally hold under either no additional conditions or for suitable constraints on $\mathcal{Y}$ which can be directly checked simply on the basis of the support of the data analyzed. From a practical perspective, this is an important gain relative to the need of verifying more sophisticated regularity conditions on the assumed model and on the unknown data generating process. Notice that the boundedness condition on $\mathcal{Y}$ in Example 3.5 interestingly relates to Assumptions 1 and 2 of Bernton et al. (2019) which have been verified when $\mathcal{Y}$ is a bounded subset of $\mathbb{R}^{d}$ by, e.g., Weed and Bach (2019). In this context, our Rademacher complexity perspective further refines the important results in Bernton et al. (2019) by clarifying that it is possible to derive convergence and concentration results for Wasserstein-abc that are regulated by a known complexity measure and hold uniformly over the space of probability measures defined on a bounded $\mathcal{Y}$.
Notice that an alternative possibility to verify Assumptions 1–2 in Bernton et al. (2019) is to leverage the results in Fournier and Guillin (2015) who replace the boundedness condition in Weed and Bach (2019) with assumptions on the existence of exponential moments; see also the supplementary materials in Bernton et al. (2019). A similar direction within our Rademacher complexity framework would be to assume that the class of functions $\mathfrak{F}$ defining the Wasserstein distance admits a uniform Glivenko–Cantelli property over a subset $\tilde{\mathcal{P}}(\mathcal{Y})$ of $\mathcal{P}(\mathcal{Y})$ that comprises probability measures on $\mathcal{Y}$ meeting some suitable regularity conditions. Recalling the previous discussion on the connection among (5) and our assumptions, this would imply a relaxation of (III)–(IV) allowing the theory in Sections 3.1 and 3.2 to still hold for statistical models and data generating processes belonging to $\tilde{\mathcal{P}}(\mathcal{Y})$. Propositions 4.3 and 4.4 explore results along these lines for mmd with unbounded kernel in $\mathbb{R}^{d}$ (which also includes abc with unbounded summary statistics), and for the Wasserstein-1 distance in $\mathbb{R}^{d}$, respectively, which do not meet (III) and (IV) when $\mathcal{Y}$ is unbounded. These two propositions clarify that concentration results can still be stated under alternative, yet related, proofs based on suitable relaxations of (III)–(IV). Nonetheless, these relaxations require checking that $\{\mu_{\theta}:\theta\in\Theta\subseteq\mathbb{R}^{p}\}$ and $\mu^{*}$ meet the conditions characterizing $\tilde{\mathcal{P}}(\mathcal{Y})$, which can be difficult since, again, $\mu^{*}$ is generally not known in practice. Conversely, when (III)–(IV) hold (e.g., in mmd with bounded kernels and Wasserstein-1 distance in bounded subsets of $\mathbb{R}^{d}$) the convergence and concentration results in Section 3.13.2 are guaranteed without the need to worry about the peculiar properties of the assumed model and of the generally-unknown data generating process.
4 Asymptotic properties of ABC with maximum mean discrepancy and Wasserstein-1 distance
Sections 4.14.2 specialize the theory derived within Section 3 to two remarkable distances in the ips class; namely mmd (which includes distances among summaries as a special case) and Wasserstein-1 distance, respectively. Recalling Examples 3.53.7 these discrepancies are covered by the general results in Section 3, under the assumption that the kernel or the space $\mathcal{Y}$ are bounded. For completeness, we extend such concentration results also to situations where these two conditions are not met; see Propositions 4.3 and 4.4.
4.1 Asymptotic properties of ABC with maximum mean discrepancy
As discussed in Sections 13, mmd stands out as a prominent example of discrepancy within summary-free abc. Nonetheless, an in-depth and comprehensive study on the limiting properties of mmd-abc is still lacking. In fact, no theory is available in the original article proposing mmd-abc methods (Park, Jitkrittum and Sejdinovic, 2016), while convergence in the fixed $\varepsilon_{n}=\varepsilon$ and $n\rightarrow\infty$ regime is explored in Jiang, Wu and Wong (2018) but without conclusive results. Nguyen et al. (2020) study convergence and concentration of abc with the energy distance in both fixed and vanishing $\varepsilon_{n}$ settings. Recalling Example 2.3, the direct correspondence between mmd and the energy distance would allow to translate these results to the mmd framework. However, as highlighted by the same authors, the theory derived relies on difficult-to-verify existence assumptions which yield bounds depending on control functions that are not made explicit.
Leveraging Theorems 3.13.3 and Corollary 3.2 along with the available upper bounds on the Rademacher complexity of mmd — see Example 3.6 — Corollaries 4.14.2 substantially refine and expand available knowledge on the limiting properties of mmd-abc with routinely-implemented bounded kernels. Crucially, mmd with such kernels automatically satisfies (III)–(IV) without additional constraints on the model or on the data generating process. Notice that Corollaries 4.14.2 also hold for summary-based distances with bounded summaries as a direct consequence of the discussion in Examples 2.4 and 3.7.
Corollary 4.1.
Consider the mmd with a bounded kernel $k(\cdot,\cdot)$ defined on $\mathbb{R}^{d}$, where $|k(x,x)|\leq 1$ for any $x\in\mathbb{R}^{d}$. Then, under (I), the acceptance probability $p_{n}$ of the rejection-based abc routine employing discrepancy $\mathcal{D}_{\textsc{mmd}}$ and the threshold $\varepsilon$ satisfies
$p_{n}\rightarrow\pi\{\theta:\mathcal{D}_{\textsc{mmd}}(\mu_{\theta},\mu^{*})\leq\varepsilon\},$
almost surely with respect to $y_{1:n}\stackrel{{\scriptstyle\mbox{\scriptsize i.i.d.}}}{{\sim}}\mu^{*}$, as $n\rightarrow\infty$, for any $\mu_{\theta},\mu^{*}\in\mathcal{P}(\mathcal{Y})$. Moreover, let $\tilde{\varepsilon}$ be defined as in Theorem 3.1. Then, for any fixed $\varepsilon>\tilde{\varepsilon}$, it holds that
$\pi_{n}^{(\varepsilon)}(\theta)\rightarrow\pi(\theta\mid\mathcal{D}_{\textsc{mmd}}(\mu_{\theta},\mu^{*})\leq\varepsilon)\propto\pi(\theta)\mathds{1}\left\{\mathcal{D}_{\textsc{mmd}}(\mu_{\theta},\mu^{*})\leq\varepsilon\right\},$
almost surely with respect to $y_{1:n}\stackrel{{\scriptstyle\mbox{\scriptsize i.i.d.}}}{{\sim}}\mu^{*}$, as $n\rightarrow\infty$.
The above result applies, e.g., to routinely-implemented Gaussian $k(x,x^{\prime})=\exp(-\|x-x^{\prime}\|^{2}/\sigma^{2})$ and Laplace $k(x,x^{\prime})=\exp(-\|x-x^{\prime}\|/\sigma)$ kernels on $\mathbb{R}^{d}$, which are both bounded by $1$, thereby implying $||f||_{\infty}\leq 1$ and $\mathfrak{R}_{n}(\mathfrak{F})\leq n^{-1/2}$. These results are also crucial to prove the concentration statement in Corollary 4.2 below.
Corollary 4.2.
Consider the mmd with a bounded kernel $k(\cdot,\cdot)$ defined on $\mathbb{R}^{d}$, where $|k(x,x)|\leq 1$ for any $x\in\mathbb{R}^{d}$. Then, for any $\mu_{\theta},\mu^{*}\in\mathcal{P}(\mathcal{Y})$, under (I)–(II) and the settings of Theorem 3.3, with $\bar{\varepsilon}_{n}=[\log(n)/n]^{1/2}$, we have that
$\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}\Bigl{(}\Bigl{\{}\theta:\mathcal{D}_{\textsc{mmd}}(\mu_{\theta},\mu^{*})>\varepsilon^{*}+\Bigl{(}\frac{10}{3}+(L+2)^{1/2}\Bigr{)}\cdot\Bigl{(}\frac{\log n}{n}\Bigr{)}^{1/2}\Bigl{\}}\Bigr{)}\leq\frac{2\cdot 3^{L}}{c_{\pi}n},$
with $\mathbb{P}_{y_{1:n}}$–probability going to $1$ as $n\rightarrow\infty$.
Notice that $\mathfrak{R}_{n}(\mathfrak{F})\leq n^{-1/2}$ implies $\mathfrak{R}_{n}(\mathfrak{F})\log\log(n)\leq\log\log(n)/n^{1/2}\leq[\log(n)/n]^{1/2}$ for any $n\geq 1$. Hence, as a consequence of the previous discussion, the concentration rate is essentially minimized by setting $\bar{\varepsilon}_{n}=[\log(n)/n]^{1/2}$. This result interestingly aligns with the optimal rate derived by Frazier et al. (2018) for summary-based abc under sub-Gaussian assumptions. The reason for such an agreement is immediately clear after noticing that mmd with bounded kernels includes, as a special case, abc with bounded summaries.
Corollaries 4.14.2 are effective examples of the potentials of Theorems 3.13.3 and Corollary 3.2, which can be readily specialized to any discrepancy within the ips class. For example, in the context of mmd with Gaussian and Laplace kernels, Corollary 4.2 ensures informative posterior concentration without requiring assumptions on $\{\mu_{\theta}:\theta\in\Theta\subseteq\mathbb{R}^{p}\}$ or $\mu^{*}$. Similar results can be obtained for all ips discrepancies as long as (III)–(IV) are satisfied and $\mathfrak{R}_{n}(\mathfrak{F})$ admits explicit upper bounds. For instance, if $\mathcal{Y}$ is bounded, this is possible for the Wasserstein-1 distance in Example 3.5, leveraging the bounds for $\mathfrak{R}_{n}(\mathfrak{F})$ in Sriperumbudur et al. (2012).
While (III) and (IV) hold for mmd with a bounded kernel without additional assumptions, the currently-available bounds on the Rademacher complexity ensure that mmd with an unbounded kernel meets the above conditions only under specific models and data generating processes, even within the i.i.d. setting. In this context, it is however possible to revisit the results for the Wasserstein case in Proposition 3 of Bernton et al. (2019) under the new Rademacher complexity framework introduced in the present article. In particular, as shown in Proposition 4.3, under mmd with an unbounded kernel, the existence Assumptions 1 and 2 in Bernton et al. (2019) can be directly related to constructive conditions on the kernel, inherently related to our Assumption (IV). This in turn yields informative concentration inequalities that are reminiscent of those in Theorem 3.3 and Corollary 4.2. Notice that these inequalities also hold for summary-based abc with routinely-used unbounded summaries (e.g., moments) as a direct consequence of the discussion in Example 3.7.
Proposition 4.3.
Consider the mmd with unbounded kernel $k(\cdot,\cdot)$ on $\mathbb{R}^{d}$. Assume (I)–(II) along with (A1) $\mathbb{E}_{y}\left[k(y,y)\right]<\infty$, (A2) $\int_{\Theta}\mathbb{E}_{z}\left[k(z,z)\right]\pi({\rm d}\theta)<\infty$, and (A3) there exist constants $\delta_{0}>0$ and $c_{0}>0$ such that $\mathbb{E}_{z}\left[k(z,z)\right]<c_{0}$ for any $\theta$ satisfying $(\mathbb{E}_{z,z^{\prime}}\left[k(z,z^{\prime})\right]-2\mathbb{E}_{z,y}\left[k(y,z)\right]+\mathbb{E}_{y,y^{\prime}}\left[k(y,y^{\prime})\right])^{1/2}\leq\varepsilon^{*}+\delta_{0}$, where $z,z^{\prime}\sim\mu_{\theta}$ and $y,y^{\prime}\sim\mu^{*}$. Then, when $n\rightarrow\infty$, $\bar{\varepsilon}_{n}\rightarrow 0$ and $n\bar{\varepsilon}^{2}_{n}\rightarrow\infty$, for some $C\in(0,\infty)$ and any $M_{n}\in(0,\infty)$, it holds that
$\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}\Bigl{(}\Bigl{\{}\theta:\mathcal{D}_{\textsc{mmd}}(\mu_{\theta},\mu^{*})>\varepsilon^{*}+\frac{4\bar{\varepsilon}_{n}}{3}+\Bigr{(}\frac{M_{n}}{n\bar{\varepsilon}_{n}^{L}}\Bigr{)}^{1/2}\Bigr{\}}\Bigr{)}\leq\frac{C}{M_{n}},$
with $\mathbb{P}_{y_{1:n}}$–probability going to $1$ as $n\rightarrow\infty$.
A popular example of unbounded kernel is provided by the polynomial one, which is defined as $k(x,x^{\prime})=(1+a\left<x,x^{\prime}\right>)^{q}$ for some integer $q\in\{2,3,\dots\}$, and constant $a>0$. Under such a kernel, it can be easily shown that if $\mathbb{E}_{y}[\|y\|^{q}]<\infty$, and $\theta\mapsto\mathbb{E}_{z}(\|z\|^{q})$ is $\pi$-integrable, then Assumptions $(A1)$$(A3)$ are satisfied. The latter conditions essentially require that the kernel has finite expectation under both $\mu^{*}$ and $\mu_{\theta}$ for suitable $\theta\in\bar{\Theta}\subseteq\Theta$, and is uniformly bounded for those $\mu_{\theta}$ close to $\mu^{*}$. Recalling Example 2.3 and the bound $\mathfrak{R}_{\mu,n}(\mathfrak{F})\leq[\mathbb{E}_{x\sim\mu}k(x,x)/n]^{1/2}$, $(A1)$$(A3)$ are inherently related to (IV), which however additionally requires that these expectations are finite for any $\mu\in\mathcal{P}(\mathcal{Y})$. Notice that while the concentration result in Proposition 4.3 does not make explicit the dependence on the regularity of the different moments (e.g., Frazier et al., 2018), as clarified in the polynomial kernel example, such a dependence on $q$ is embedded within the definition of the kernel, which in turn enters the expression for $\mathcal{D}_{\textsc{mmd}}$. As such, translating the results in Proposition 4.3 from our overarching focus on the space of distributions to the one of parameters would make this dependence explicit. Such a mapping is challenging to obtain analytically in general settings. Hence, further refinements along these lines are left for future research.
Before moving to the Wasserstein-1 distance, notice that by Proposition 4.3 a sensible setting for $\bar{\varepsilon}_{n}$ in the unbounded-kernel case would be $\bar{\varepsilon}_{n}=(M_{n}/n)^{1/(2+L)}$. This yields
${\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}}(\{\theta:\mathcal{D}_{\textsc{mmd}}(\mu_{\theta},\mu^{*})>\varepsilon^{*}+(7/3)(M_{n}/n)^{1/(2+L)}\})\leq C/M_{n},$
which is essentially the tighter possible order of magnitude for the bound. In the unbounded-kernel setting, $M_{n}=n$ would not be suitable, but any $M_{n}\rightarrow\infty$ slower than $n$ can work; e.g., $M_{n}=n^{1/2}$ yields $\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}(\{\theta:\mathcal{D}_{\textsc{mmd}}(\mu_{\theta},\mu^{*})>\varepsilon^{*}+(7/3)(1/n)^{1/(2L+4)}\})\leq C/n^{1/2}.$
4.2 Asymptotic properties of abc with Wasserstein-1 distance
As pointed out earlier, the general results in Section 3 apply directly to the Wasserstein-1 distance ($\mathcal{D}_{\textsc{wass}}$) when the data space $\mathcal{Y}$ is bounded. As clarified in Proposition 4.4, if $\mathcal{Y}$ is unbounded, concentration results can still be derived for $\mathcal{D}_{\textsc{wass}}$, but at the expense of regularity conditions on $\{\mu_{\theta}:\theta\in\Theta\subseteq\mathbb{R}^{p}\}$ and $\mu^{*}$. These results and the associated proof mirror those for the mmd with unbounded kernel in Proposition 4.3, and rely on exponential moment assumptions that allow to leverage the recent concentration bounds for $\mathcal{D}_{\textsc{wass}}$ in Lei (2020).
Proposition 4.4.
Consider the Wasserstein-1 distance $\mathcal{D}_{\textsc{wass}}$ (see Example 2.2) on $\mathcal{Y}=\mathbb{R}^{d}$, with ground distance $\rho(x,x^{\prime})=\|x-x^{\prime}\|$. Besides (I)–(II), assume there exists a small enough $c>0$ such that (A1’) $\ \mathbb{E}_{y}(\exp(c\|y\|))<\infty$ and (A2’) ${\sup_{\mu_{\theta},\theta\in\Theta}}\mathbb{E}_{z}(\exp(c\|z\|))<~{}\infty$, where $z\sim\mu_{\theta}$ and $y\sim\mu^{*}$. Then, when $n\rightarrow\infty$ and $\bar{\varepsilon}_{n}\rightarrow 0$ such that $n\bar{\varepsilon}^{2}_{n}\rightarrow\infty$ and also $n^{-1/\max(d,3)}\ll\bar{\varepsilon}_{n}$, for some $C\in(0,\infty)$ and any $M_{n}\in(0,\infty)$, it holds that
$\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}\Bigl{(}\Bigl{\{}\theta:\mathcal{D}_{\textsc{wass}}(\mu_{\theta},\mu^{*})>\varepsilon^{*}+\frac{4\bar{\varepsilon}_{n}}{3}+\Bigl{(}\frac{1}{c^{\prime}n}\log\frac{M_{n}}{\bar{\varepsilon}_{n}^{L}}\Bigr{)}^{1/2}+c_{1}n^{-{1}/{\max(d,3)}}\Bigr{\}}\Bigr{)}\leq\frac{C}{M_{n}},$
with $\mathbb{P}_{y_{1:n}}$–probability going to $1$ as $n\rightarrow\infty$.
The results and proof of Proposition 4.4 are inherently related to those in Section 2.1 of the supplementary materials in Bernton et al. (2019). However, rather than leveraging the concentrations bounds by Fournier and Guillin (2015), Proposition 4.4 exploits the more recent ones in Lei (2020) under the exponential moment assumptions in (A1’)–(A2’) that are generally satisfied for sub-exponential random variables, including sub-Gaussian ones. We refer to Lei (2020) for a more detailed discussion on the advantages of these bounds compared to those in Fournier and Guillin (2015). Notice that, under Proposition 4.4, setting $\bar{\varepsilon}_{n}=n^{-1/\max(d,3)}\sqrt{\log(n)}$ is an admissible choice to ensure concentration, but the induced rate would be worse than the one in Theorem 3.3. Deriving optimal rates for the Wasserstein-1 distance in unbounded $\mathcal{Y}$ is not the scope of Proposition 4.4, whose aim is, instead, to clarify that concentration beyond the settings considered in Section 3 can still be achieved, but at the cost of regularity conditions on $\{\mu_{\theta}:\theta\in\Theta\subseteq\mathbb{R}^{p}\}$ and $\mu^{*}$ that do not allow to state results uniformly over $\mathcal{P}(\mathcal{Y})$. As such, we leave questions on the optimality of the bounds for the Wasserstein distance in unbounded spaces $\mathcal{Y}$ as future research, and refer to Bernton et al. (2019) for more in-depth and specialized results on the properties of abc under such a distance.
5 Illustrative simulation in i.i.d. settings
Let us illustrate the theory we derived in the previous sections through a simple, yet insightful, simulation in i.i.d. settings (see Appendix C in the Supplementary Material for empirical results under non-i.i.d. regimes). Note that several empirical studies have already compared the performance of summary-free abc under different discrepancies and complex non-i.i.d. models. Recalling Drovandi and Frazier (2022), all these analyses clarify the practical feasibility of abc based on various discrepancies, including those within the ips class, and in several i.i.d. and non-i.i.d. scenarios which require abc procedures. This feasibility is also supported by recent softwares (e.g., Dutta et al., 2021).
Rather than replicating the available studies on benchmark examples, we complement current empirical evidence on summary-free abc by focusing on a misspecified and contaminated scenario that clarifies the possible challenges in convergence and concentration encountered even in basic i.i.d. settings. As clarified in Table 1 and Figure 1, this scenario also showcases a key consequence of the novel theoretical results in Sections 24. Namely that effective ips discrepancies with guarantees of uniform convergence and concentration are a safe and sensibile choice in the absence of knowledge on the specific properties of $\mu^{*}$.
Table 1: Concentration and runtimes in seconds (for a single discrepancy evaluation) of abc under mmd with Gaussian kernel, Wasserstein-1 distance, summary-based distance (mean) and kl divergence for a misspecified Huber contamination model with a varying $\alpha\in\{0.05,0.10,0.15\}$. $\textsc{mse}=\hat{\mathbb{E}}_{\mu^{*}}[\hat{\mathbb{E}}_{\textsc{abc}}(\theta-\theta_{0})^{2}]$.
mse ($\alpha=0.05$) mse ($\alpha=0.10$) mse ($\alpha=0.15$) time
(ips) mmd 0.024 0.027 0.031 $<$ 0.01”
(ips) Wasserstein-1 0.027 0.067 0.122 $<$ 0.01”
(ips) summary (mean) 0.841 2.648 2.835 $<$ 0.01”
(non-ips) kl 0.073 0.076 0.077 $<$ 0.01”
Recalling Example 3.4, we consider, in particular, an uncontaminated bivariate Student’s $t$ distribution $\mu_{\theta_{0}}$ with $3$ degrees of freedom, mean vector $(1,1)$, and dispersion matrix having entries $\sigma_{11}=\sigma_{22}=1$ and $\sigma_{12}=\sigma_{21}=0.5.$ Such an uncontaminated data generating process is then perturbed with three different levels $\alpha_{n}=\alpha\in\{0.05,0.10,0.15\}$ of contamination from a Student’s $t$ distribution $\mu_{C}$ having the same parameters as $\mu_{\theta_{0}}$, except for the mean vector which is set to $(20,20)$. As such, the data $y_{1:n}$ from $\mu^{*}=(1-\alpha)\mu_{\theta_{0}}+\alpha\mu_{C}$ are obtained by sampling $n=100$ draws from the bivariate Student’s $t$ $\mu_{\theta_{0}}$ and then replacing $(100\cdot\alpha)\%$ of these draws with samples from the contaminating Student’s $t$ distribution $\mu_{C}$. For Bayesian inference, we focus on the parameter $\theta\in\mathbb{R}$ defining the unknown location vector $[1,1]^{\intercal}\theta$, and consider a misspecified bivariate Gaussian model $\mu_{\theta}$ with mean vector $[1,1]^{\intercal}\theta$ and known covariance matrix coinciding with that of the uncontaminated Student’s $t$ data generating process. Such a choice is interesting in providing a model that is slightly misspecified even when the data are not contaminated. Notice that, although this model does not necessarily require an abc approach to allow Bayesian inference, as discussed above, the issues outlined in Table 1 and Figure 1 for certain discrepancies, even in such a basic example, provide a useful empirical insight that complements those in the extensive quantitive studies already available in the literature for more complex settings.
Figure 1: Graphical representation of the abc posterior for $\theta$ under mmd with Gaussian kernel, Wasserstein-1 distance, summary-based distance (mean) and kl divergence for one simulated dataset from a misspecified Huber contamination model with a varying $\alpha\in\{0.05,0.10,0.15\}$; see also Example 3.4 for details. The red dashed line corresponds to the location parameter $\theta_{0}=1$ of the uncontaminated model.
In performing abc under the above model and for the different discrepancies of interest, we employ rejection-based abc with $m=n=100$ and a $\mbox{N}(0,1)$ prior for $\theta$. Following the standard practice in comparing discrepancies (Drovandi and Frazier, 2022), we specify a common budget of $T=25{,}000$ simulations and define the abc threshold to retain, for every discrepancy analyzed, the $1\%$ (i.e., $250$) values of $\theta$ that generated the synthetic data closest to the observed ones, under such a discrepancy. Although the theory in Section 3 can potentially guide the choice of the threshold, similar guidelines are not yet available beyond the ips class. Hence, to ensure a fair assessment across all the discrepancies, we rely on the recommended practice in comparing the different abc implementations. As clarified in Table 1 and Figure 1, the discrepancies assessed are those most commonly used in abc, namely, the Wasserstein-1 distance as in Example 2.2 (Bernton et al., 2019), mmd with Gaussian kernel defined in Example 2.3 (Park, Jitkrittum and Sejdinovic, 2016; Nguyen et al., 2020), and a summary-based distance leveraging the sample mean as the summary statistics (see Example 2.4). For comparison, we also consider a popular discrepancy that does not belong to the ips class, i.e., the Kullback–Leibler divergence (Jiang, Wu and Wong, 2018). All these discrepancies can be effectively implemented in R, leveraging, in particular, the libraries transport and eummd. For mmd, the choice of the length-scale parameter $\sigma^{2}$ is based on the median heuristic (Gretton et al., 2012) automatically implemented in the function mmd.
Leveraging the samples from the abc posterior for $\theta$ under each discrepancy, we first estimate the mean squared error $\hat{\mathbb{E}}_{\textsc{abc}}(\theta-\theta_{0})^{2}$ with respect to the location $\theta_{0}=1$ of the uncontaminated Student’s $t$, under the different discrepancies, thereby assessing performance with a focus on a common metric on the space of parameters. Table 1 displays these error estimates, under each discrepancy and level of contamination, further averaged over $50$ simulated datasets to obtain a comprehensive assessment based on replicated studies. The results in Table 1 together with the graphical representation in Figure 1 of the abc posteriors for one simulated dataset from the contaminated model with varying $\alpha\in\{0.05,0.10,0.15\}$ effectively illustrate an important practical consequence of the theory derived in Sections 24. Namely, when one is not sure, or cannot check, whether the assumed statistical model and/or the underlying data generating process meet specific regularity conditions, effective ips discrepancies with guarantees of uniform convergence and concentration (e.g., mmd with bounded kernel) provide a robust and safe default choice. When the level of contamination is mild ($\alpha=0.05$), Wasserstein-abc achieves comparable concentration. This result aligns with Proposition 4.4, which however depends on the specific properties of the model and the underlying data generating process. Such a dependence is evident when the amount of contamination grows to $\alpha=0.10$ and $\alpha=0.15$. In these settings, the performance of Wasserstein-abc tends to slightly deteriorate, mainly due to a location shift in the induced posterior. This shift is even more evident for summary-based abc relying on the sample mean, that is not robust to location contaminations. Conversely, the Kullback–Leibler divergence preserves robustness but at the expense of a lower concentration. Note that all the discrepancies analyzed in Table 1 and Figure 1 are well-defined under the Student’s $t$ and Gaussian distributions considered in this simulation study. In particular, since the Student’s $t$ has $3$ degrees of freedom, its mean and variance are finite. Thus, Wasserstein-1 is well-defined for both the assumed model and the underlying data generating process.
Considering the running times, as displayed in Table 1 all the discrepancies under analysis can be evaluated in the order of milliseconds for a sample of size $n=m=100$ on a standard laptop. This enables scalable and effective R implementations.
6 Discussion
This article provides theoretical advancements with respect to the recent literature on the asymptotic properties of discrepancy-based abc posteriors by connecting these properties with the behavior of the Rademacher complexity associated with the chosen discrepancy. As clarified in the article and in the Supplementary Material, although the Rademacher complexity has never been considered in abc, this notion yields a powerful and promising perspective to derive general, informative and uniform convergence and concentration properties.
While the above contribution already provides key advancements, the proposed perspective based on Rademacher complexity has broader scope and sets the premises for additional future research. For example, as clarified in this article, any novel result and bound on the Rademacher complexity of specific discrepancies can be directly applied to abc theory through our framework. This may yield tighter or more explicit bounds, possibly holding under milder assumptions and more general discrepancies. For instance, to our knowledge, informative bounds for the Rademacher complexity of the Wasserstein-1 distance are currently available only for bounded $\mathcal{Y}$ and, hence, it would be of interest to leverage future findings on the unbounded $\mathcal{Y}$ case to further refine Proposition 4.4 and broaden the range of models for which our theory, when specialized to Wasserstein-1 distance, applies. To this end, it might also be promising to explore results on local Rademacher complexities (e.g., Bartlett, Bousquet and Mendelson, 2005), along with the proposed variable transformation strategy in Example 3.5. This solution requires the choice of a mapping $g(\cdot)$ which clearly influences the learning properties of the induced distance and, as such, requires further investigation. Finally, although Appendix C in the Supplementary Material provides important extensions to non-i.i.d. settings, other relaxations beyond $\beta$-mixing processes could be of interest, such as, for instance, the case of independent but not identically distributed data. This setting can be addressed via the residual reconstruction strategy (e.g., Bernton et al., 2019, Section 4.2.3.) that would imply studying discrepancies among empirical distributions of residuals, for which i.i.d. assumptions are again reasonable.
Although our focus is on the active convergence and concentration theory for discrepancy-based abc, other properties such as accuracy in uncertainty quantification of credible intervals and the limiting shapes of abc posteriors, in correctly-specified models, have attracted interest in the context of summary-based and summary-free abc (Frazier et al., 2018; Frazier, 2020; Wang, Kaji and Rockova, 2022). While this direction goes beyond the scope of our article, extending and unifying such results, as done for concentration properties, is also of interest.
While the ips class is broad, it does not cover all discrepancies employed in abc. For example, the kl divergence (Jiang, Wu and Wong, 2018) and the Hellinger distance (Frazier, 2020) are not ips, but rather belong to the class of $f$-divergences. While this latter family has important differences relative to the ips class, it is worth investigating $f$-divergences in the light of our results. To accomplish this goal, an option is to exploit the unified treatment of these two classes in, e.g., Agrawal and Horel (2021) and Birrell et al. (2022). More generally, our results could also stimulate methodological and theoretical advancements in generalized likelihood-free Bayesian inference via discrepancy-based pseudo-posteriors (e.g., Bissiri, Holmes and Walker, 2016; Jewson, Smith and Holmes, 2018; Miller and Dunson, 2019; Chérief-Abdellatif and Alquier, 2020; Matsubara et al., 2022). The recent contribution by Frazier, Knoblauch and Drovandi (2024) establishes, among other key results, important connections between such a latter framework and abc that could facilitate these advancements.
Finally, notice that, for the sake of simplicity and ease of comparison with related studies, we have focused on rejection abc, and constrained the number $m$ of synthetic samples to be equal to the sample size $n$ of the observed data. While these settings are standard in state-of-the-art theory (Bernton et al., 2019; Frazier, 2020), other abc routines and alternative scenarios where $m$ grows, e.g., sub-linearly, with $n$ deserve further investigation. This latter regime would be of interest in settings where the simulation of synthetic data is computationally expensive.
{acks}
[Acknowledgments] We are grateful to the Editor, the Associate Editor and the referees for the constructive feedbacks, which helped us in improving the preliminary version of the article.
[…]
14pt
SUPPLEMENTARY MATERIAL TO “CONCENTRATION OF DISCREPANCY-BASED APPROXIMATE BAYESIAN
COMPUTATION VIA RADEMACHER COMPLEXITY”
Appendix A Additional Integral Probability Semimetrics
A.1 Two additional examples of ips discrepancies
While mmd, Wasserstein-1 and summary-based distances provide the most notable examples of ips discrepancies employed in abc, two other relevant ips instances are the total variation (tv) distance and the Kolmogorov–Smirnov (ks) distance, discussed below.
Example A.1.
(Total variation distance). Although the total variation distance is not a common choice within discrepancy-based abc, it still provides a notable example of ips, obtained when $\mathfrak{F}$ is the class of measurable functions whose sup-norm is bounded by $1$; i.e. $\mathfrak{F}=\{f:||f||_{\infty}\leq 1\}$.
Example A.2.
(Kolmogorov–Smirnov distance). When $\mathcal{Y}=\mathbb{R}$ and $\mathfrak{F}=\{\mathds{1}_{(-\infty,a]}\}_{a\in\mathbb{R}}$, then $\mathcal{D}_{\mathfrak{F}}$ is the Kolmogorov–Smirnov distance, which can also be written as $\mathcal{D}_{\mathfrak{F}}(\mu_{1},\mu_{2})=\sup_{x\in\mathcal{Y}}|F_{1}(x)-F_{2}(x)|$, where $F_{1}$ and $F_{2}$ are the cumulative distribution functions associated with $\mu_{1}$ and $\mu_{2}$, respectively.
A.2 Validity of (III)–(IV) for the tv distance and Kolmogorov–Smirnov distance
Examples A.3 and A.4 verify the validity of assumptions (III) and (IV) under the tv distance and the Kolmogorov–Smirnov distance, respectively.
Example A.3.
(Total variation distance). The tv distance satisfies (III) by definition, but in general not Assumption (IV), unless the cardinality $|\mathcal{Y}|$ of $\mathcal{Y}$ is finite. In fact, when $\mathcal{Y}=\mathbb{R}$ and $\mu\in\mathcal{P}(\mathcal{Y})$ is continuous, the probability that there exists an index $i\neq i^{\prime}$ such that $x_{i}=x_{i^{\prime}}$ is zero. Hence, with probability $1$, for any vector $\epsilon_{1:n}$ of Rademacher variables there always exists a function $f_{\epsilon}$ from $\mathcal{Y}$ to $\{0;1\}$ such that $f_{\epsilon}(x_{i})=\mathds{1}_{\{\epsilon_{i}=1\}}$. Therefore, $\sup_{f\in\mathfrak{F}}|(1/n)\sum_{i=1}^{n}\epsilon_{i}f(x_{i})|\geq(1/n)\sum_{i=1}^{n}\mathds{1}_{\{\epsilon_{i}=1\}}$, which implies that the Rademacher complexity $\mathfrak{R}_{\mu,n}(\mathfrak{F})$ is bounded below by $(1/n)\sum_{i=1}^{n}\mathbb{P}(\epsilon_{i}=1)=1/2.$ Nonetheless, as mentioned above, the tv distance can still satisfy (IV) in specific contexts. For instance, leveraging the bound in Lemma 5.2 of Massart (2000), when the cardinality $|\mathcal{Y}|$ of $\mathcal{Y}$ is finite, there will be replicates in $[f(x_{1}),\ldots,f(x_{n})]$ whenever $n>|\mathcal{Y}|$. Hence, as $n\rightarrow\infty$, it will be impossible to find a function in $\mathfrak{F}$ which can interpolate any noise vector of Rademacher variables with $[f(x_{1}),\ldots,f(x_{n})]$, thus ensuring $\mathfrak{R}_{n}(\mathfrak{F})\rightarrow 0$.
Example A.4.
(Kolmogorov–Smirnov distance). The ks distance meets (III) by definition and, similarly to mmd with bounded kernels, also condition (IV) is satisfied without the need to impose additional constraints on the model $\mu_{\theta}$ or on the data-generating process. More specifically, Assumption (IV) follows from the inequality $\mathfrak{R}_{\mu,n}(\mathfrak{F})\leq 2[\log(n+1)/n]^{1/2}$ in Chapter 4.3.1 of Wainwright (2019). This is a consequence of the bounds on $\mathfrak{R}_{\mu,n}(\mathfrak{F})$ when $\mathfrak{F}$ is a class of $b$-uniformly bounded functions such that, for some $\nu\geq 1$, it holds $\mbox{card}\{f(x_{1:n}):f\in\mathfrak{F}\}\leq(n+1)^{\nu}$ for any $n$ and $x_{1:n}$ in $\mathcal{Y}^{n}$. When $\mathfrak{F}=\{1_{(-\infty,a]}\}_{a\in\mathbb{R}}$ each $x_{1:n}$ would divide the real line in at most $n+1$ intervals and every indicator function within $\mathfrak{F}$ will take value $1$ for all $x_{i}\leq a$ and zero otherwise, meaning that $\mbox{card}\{f(x_{1:n}):f\in\mathfrak{F}\}\leq(n+1)$. Therefore, by applying Equation (4.24) in Wainwright (2019), with $b=1$ and $\nu=1$, yields $\mathfrak{R}_{\mu,n}(\mathfrak{F})\leq 2[\log(n+1)/n]^{1/2}$ for any $\mu\in\mathcal{P}(\mathcal{Y})$, which implies that Assumption (IV) is met. These derivations clarify the usefulness of the available techniques for upper bounding the Rademacher complexity (e.g., Wainwright, 2019, Chapter 4.3), leveraging, in this case, the notion of polynomial discrimination and the closely-related vc dimension.
Appendix B Concentration in the space of parameters
Theorem 3.3 in the main article is stated for neighborhoods within the space of distributions. Although such a perspective is in line with the overarching focus of current theory for discrepancy-based abc (Jiang, Wu and Wong, 2018; Bernton et al., 2019; Nguyen et al., 2020; Frazier, 2020; Fujisawa et al., 2021), it shall be emphasized that similar results can be also derived in the space of parameters. To this end, it suffices to adapt Corollary 1 in Bernton et al. (2019) to our general framework, under the same additional assumptions, which are adapted below to the whole ips class.
  1. (V)
    The minimizer $\theta^{*}$ of $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ exists and is well separated, meaning that for any $\delta>0$ there is a $\delta^{\prime}>0$ such that $\inf_{\theta\in\Theta:d(\theta,\theta^{*})>\delta}\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})>\mathcal{D}_{\mathfrak{F}}(\mu_{\theta^{*}},\mu^{*})+\delta^{\prime}$;
  2. (VI)
    The parameters $\theta$ are identifiable, and there exist positive constants $K>0$, $\nu>0$ and an open neighborhood $U\subset\Theta$ of $\theta^{*}$ such that, for any $\theta\in U$, it holds that $d(\theta,\theta^{*})\leq K\left[\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})-\varepsilon^{*}\right]^{\nu}.$
Assumptions (V) and (VI) essentially require that the parameters $\theta$ are identifiable, sufficiently well-separated, and that the distance $d(\cdot,\cdot)$ between parameter values has some reasonable correspondence with the discrepancy $\mathcal{D}_{\mathfrak{F}}(\cdot,\cdot)$ among the associated distributions. Although these two assumptions introduce a condition on the model, it shall be emphasized that (V) and (VI) are not specific to our framework (e.g., Frazier et al., 2018; Bernton et al., 2019; Frazier, 2020). On the contrary, these identifiability conditions are arguably customary and minimal requirements in parameter inference. Moreover, these two assumptions have been checked in Chérief-Abdellatif and Alquier (2022) for mmd and in Bernton et al. (2019) for Wasserstein distance, which are arguably the two most remarkable examples of ips employed in the abc context. Under (V) and (VI), it is possible to state Corollary B.1.
Corollary B.1.
Assume (I)–(IV) along with (V)–(VI), and that $\mathcal{D}_{\mathfrak{F}}$ denotes a discrepancy within the ips class in Definition 2.1. Moreover, take $\bar{\varepsilon}_{n}\to 0$ as $n\to\infty$, with $n\bar{\varepsilon}^{2}_{n}\to\infty$ and $\bar{\varepsilon}_{n}/\mathfrak{R}_{n}(\mathfrak{F})\to\infty$. Then, the abc posterior with threshold $\varepsilon_{n}=\varepsilon^{*}+\bar{\varepsilon}_{n}$ satisfies
$\smash{\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}}\Bigl{(}\Bigl{\{}\theta:d(\theta,\theta^{*})>K\Bigl{[}\frac{4\bar{\varepsilon}_{n}}{3}+2\mathfrak{R}_{n}(\mathfrak{F})+\Bigl{(}\frac{2b^{2}}{n}\log\frac{n}{\bar{\varepsilon}_{n}^{L}}\Bigr{)}^{1/2}\Bigr{]}^{\nu}\Bigr{\}}\Bigr{)}\leq\frac{2\cdot 3^{L}}{c_{\pi}n},$
with $\mathbb{P}_{y_{1:n}}$–probability going to $1$ as $n\rightarrow\infty$.
As for Theorem 3.3, also Corollary B.1 holds more generally when replacing both $n/\bar{\varepsilon}_{n}^{L}$ and $c_{\pi}n$ with $M_{n}/\bar{\varepsilon}_{n}^{L}$ and $c_{\pi}M_{n}$, respectively, for any $M_{n}>1$. The proof of Corollary B.1 follows directly from Theorem 3.3 and Assumptions (V)–(VI), thereby allowing to inherit the discussion after Theorem 3.3, also when the concentration is measured directly within the parameter space via $d(\theta,\theta^{*})$. For instance, when $d(\cdot,\cdot)$ is the Euclidean distance and $\nu=1$, this implies that whenever $\mathfrak{R}_{n}(\mathfrak{F})=\mathcal{O}(n^{-1/2})$ the contraction rate will be in the order of $\mathcal{O}([\log(n)/n]^{1/2})$, which is the expected rate in parametric models.
Appendix C Extension to non-i.i.d. settings
Although the theoretical results in Sections 24 provide an improved understanding of the limiting properties of discrepancy-based abc posteriors, the i.i.d. assumption in (I) rules out important settings which often require abc. A remarkable case is that of time-dependent observations (e.g., Fearnhead and Prangle, 2012; Bernton et al., 2019; Nguyen et al., 2020; Drovandi and Frazier, 2022).
Section C.1 clarifies that the theory derived under i.i.d.  assumptions in Section 24 can be naturally extended to these non-i.i.d. settings leveraging results for Rademacher complexity in $\beta$-mixing stochastic processes (Mohri and Rostamizadeh, 2008). Examples C.3C.4 below show that such a class embraces several processes of direct practical interest. Extensions beyond this class, albeit relevant, are challenging even when the focus is on proving simpler, non-uniform, concentration results for a single discrepancy. Hence, these extensions are left for future research, that could be facilitated by the derivation of Rademacher complexity bounds for general processes beyond the $\beta$-mixing ones studied in Mohri and Rostamizadeh (2008).
C.1 Convergence and concentration beyond i.i.d. settings
Let us assume again that $\mathcal{Y}$ is a metric space endowed with distance $\rho$. However, unlike the i.i.d. setting considered in Section 2, we now focus on the situation in which the observed data $y_{1:n}=(y_{1},\ldots,y_{n})\in\mathcal{Y}^{n}$ are dependent and drawn from the joint distribution $\mu^{*(n)}\in\mathcal{P}(\mathcal{Y}^{n})$, where $\mathcal{P}(\mathcal{Y}^{n})$ is the space of probability measures on $\mathcal{Y}^{n}$. Under this more general framework, the i.i.d. case is recovered by assuming that $\mu^{*(n)}$ can be expressed as a product, i.e., $\mu^{*(n)}=\prod_{i=1}^{n}\mu^{*}$.
In the following, the above product structure is not imposed. Instead, we only assume that the marginal of $\mu^{*(n)}$ is constant, and denoted with $\mu^{*}$. Such an assumption is met whenever $y_{1:n}$ is extracted from a stationary stochastic process $(y_{t})_{t\in\mathbb{Z}}$, thus embracing a broader variety of applications of direct interest. Under these settings, a statistical model is defined as a collection of distributions in $\mathcal{P}(\mathcal{Y}^{n})$, i.e., $\{\mu_{\theta}^{(n)}:\theta\in\Theta\subseteq\mathbb{R}^{p}\}$, with a constant marginal denoted by $\mu_{\theta}$. Notice that these assumptions of constant marginals $\mu^{*}$ and $\mu_{\theta}$ are made also in the available concentration theory under non-i.i.d. settings (see e.g., Bernton et al., 2019; Nguyen et al., 2020) when requiring convergence of $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$ and suitable concentration inequalities for $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})$. As a result, the settings we consider are not more restrictive than those addressed in discrepancy-specific theory. In fact, both Bernton et al. (2019) and Nguyen et al. (2020) explicitly refer to stationary processes when discussing the validity of the assumptions on $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$ and $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})$ in non-i.i.d. contexts.
Given the above statistical model, a prior $\pi$ on $\theta$ and a generic ips discrepancy $\mathcal{D}_{\mathfrak{F}}$, the abc posterior with threshold $\varepsilon_{n}\geq 0$ is defined as
$\pi_{n}^{(\varepsilon_{n})}(\theta)\propto\pi(\theta)\int_{\mathcal{Y}^{n}}\mathds{1}\{\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon_{n}\}\ \mu_{\theta}^{(n)}(dz_{1:n}).$
This definition is the same as the one in Section 2, with the only difference that $\mu_{\theta}^{n}=\prod_{i=1}^{n}\mu_{\theta}$ is replaced by the joint $\mu^{(n)}_{\theta}$, since now the data are no more assumed to be independent.
In order to extend the convergence result in Corollary 3.2 together with the concentration statement in Theorem 3.3 to the above framework, we require an analog of Equation (2) in Lemma 2.6 for time-dependent data. This generalization can be derived leveraging results in Mohri and Rostamizadeh (2008) under the notion of $\beta$-mixing coefficients.
Definition C.1 ($\beta$-mixing).
Consider the stationary sequence $(x_{t})_{t\in\mathbb{Z}}$ of random variables, and let $\sigma^{j^{\prime}}_{j}$ be the $\sigma$–algebra generated by the random variables $x_{k}$, $j\leq k\leq j^{\prime}$, for any $j,j^{\prime}\in\mathbb{Z}\cup\{-\infty,+\infty\}$. Then, for any integer $k>0$, the $\beta$-mixing coefficient of $(x_{t})_{t\in\mathbb{Z}}$ is defined as
$\beta(k)=\mbox{sup}_{t\in\mathbb{Z}}\mathbb{E}[\mbox{sup}_{A\in\sigma_{t+k}^{\infty}}|\mathbb{P}(A\mid\sigma_{-\infty}^{t})-\mathbb{P}(A)|].$
If $\beta(k)\to 0$ as $k\to\infty$, then the stochastic process $(x_{t})_{t\in\mathbb{Z}}$ is said to be $\beta$-mixing.
Intuitively, $\beta(k)$ measures the dependence between the past (before $t$) and the future (after $t+k$) of the process. When such a dependence is weak, we expect that $\beta(k)$ will decay to 00 fast when $k\rightarrow\infty$. In the most extreme case, when the $x_{t}$’s are i.i.d., we have $\beta(k)=0$ for all $k>0$. More generally, as clarified in Definition C.1, a process having $\beta(k)\rightarrow 0$ when $k\rightarrow\infty$ is named $\beta$-mixing. We refer the reader to Doukhan (1994) for an in-depth study of the main properties of $\beta$-mixing processes along with a more comprehensive discussion of relevant examples. The most remarkable ones will be also presented in the following.
Leveraging the notion of $\beta$-mixing coefficient, Lemma C.2 extends Lemma 2.6 to the dependent setting. The proof can be found in Appendix D and combines Proposition 2 and Lemma 2 in Mohri and Rostamizadeh (2008). For readability, let us also introduce the notation $s_{n}=\lfloor n/(2\lfloor\sqrt{n}\rfloor)\rfloor$. Note that $s_{n}\sim\sqrt{n}/2$ as $n\to\infty$, and thus $s_{n}\to\infty$.
Lemma C.2.
Define $s_{n}=\lfloor n/(2\lfloor\sqrt{n}\rfloor)\rfloor$. Moreover, consider the stationary stochastic process $(x_{t})_{t\in\mathbb{Z}}$ and denote with $\beta(k),k\in\mathbb{N}$, its $\beta$-mixing coefficients. Let $\mu^{(n)}$ be the joint distribution of a sample $x_{1:n}$ extracted from $(x_{t})_{t\in\mathbb{Z}}$ and denote with $\mu=\mu^{(1)}$ its constant marginal. Then, for any $b$-uniformly bounded class $\mathfrak{F}$, any integer $n\geq 1$ and scalar $\delta\geq 0$,
$\qquad\ \ \ \mathbb{P}_{x_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\leq 2\mathfrak{R}_{\mu,s_{n}}(\mathfrak{F})+\frac{4b}{\sqrt{n}}+\delta\right]\geq 1-2{\cdot}{\exp}\left[-\frac{s_{n}\delta^{2}}{2b^{2}}\right]-2s_{n}\beta({\lfloor}\sqrt{n}{\rfloor}),$ (C.1)
with $\mathfrak{R}_{\mu,s_{n}}(\mathfrak{F})$ the Rademacher complexity in Definition 2.5 for an i.i.d. sample of size $s_{n}$ from $\mu$.
Equation (C.1) extends (2) beyond i.i.d. settings. This extension provides a bound that still depends on the Rademacher complexity in Definition 2.5 for an i.i.d. sample — in this case from the common marginal $\mu$ of the process $(x_{t})_{t\in\mathbb{Z}}$. As such, Assumption (IV) requires no modifications, and no additional validity checks relative to those discussed in Section 3.3. This suggests that the Rademacher complexity framework might also be leveraged to derive improved convergence and concentration results for discrepancy-based abc posteriors in more general situations which do not necessarily meet Assumption (I). To prove these results we leverage again Assumptions (II), (III) and (IV), and replace (I) with condition (VII).
  1. (VII)
    The data $y_{1:n}$ are from a $\beta$-mixing stochastic process $(y_{t})_{t\in\mathbb{Z}}$ with mixing coefficients $\beta(k)\leq C_{\beta}e^{-\gamma k^{\xi}}$ for some $C_{\beta},\gamma,\xi>0$, common marginal $\mu$, and generic joint $\mu^{*(n^{\prime})}$ for a sample $y_{1:n^{\prime}}$ from $(y_{t})_{t\in\mathbb{Z}}$ for any $n^{\prime}\in\mathbb{N}$. The same $\beta$-mixing conditions hold also for the process $(z_{t})_{t\in\mathbb{Z}}$ associated with the synthetic data $z_{1:n}$ from the assumed model. In this case, the joint distribution for a generic sample $z_{1:n^{\prime}}$ is $\mu^{(n^{\prime})}_{\theta}$$\theta\in\Theta$, and the common marginal is denoted by $\mu_{\theta}$. For simplicity and without loss of generality, we also assume that the constants $C_{\beta}$, $\gamma$, and $\xi$ are the same for $(y_{t})_{t\in\mathbb{Z}}$ and $(z_{t})_{t\in\mathbb{Z}}$.
Assumption (VII) is clearly more general than (I). As discussed previously, it embraces several stochastic processes of substantial interest in practical applications, including those in Examples C.3C.4 below; see Doukhan (1994) for additional examples and discussion.
Example C.3 (Doeblin-recurrent Markov chains).
Let $(x_{t})_{t\in\mathbb{Z}}$ be a Markov chain on $\mathcal{Y}\subset\mathbb{R}^{d}$ with transition kernel $P(\cdot,\cdot)$. Such a Markov chain is said to be Doeblin-recurrent if there exists a probability measure $q$, a constant $0<c\leq 1$ and an integer $r>0$ such that, for any measurable set $A$ and any $x\in\mathbb{R}^{d}$, $P^{r}(x,A)\geq cq(A)$. When this is the case, $(x_{t})_{t\in\mathbb{Z}}$ is $\beta$-mixing with $\beta(k)\leq 2(1-c)^{k/r}$; see e.g., Theorem 1 in page 88 of Doukhan (1994).
Example C.4 (Hidden Markov chains).
Assume $(x_{t})_{t\in\mathbb{Z}}$ is a $\beta$-mixing stochastic process with coefficients $\beta_{x}(k),k\in\mathbb{N}$. If $\tilde{x}_{t}=F(x_{t},\varepsilon_{t})$ with $\varepsilon_{t}$ i.i.d., then the $\beta$-mixing coefficients of $(\tilde{x}_{t})_{t\in\mathbb{Z}}$ satisfy $\beta_{\tilde{x}}(k)=\beta_{x}(k)$. Therefore, $(\tilde{x}_{t})_{t\in\mathbb{Z}}$ is also $\beta$-mixing and inherits the bounds on $\beta_{x}(k)$. These processes are often used in practice with $(x_{t})_{t\in\mathbb{Z}}$ being a Markov chain. In this case $(\tilde{x}_{t})_{t\in\mathbb{Z}}$ is called a Hidden Markov chain.
Section 2.4.2 of Doukhan (1994) also provides conditions on $F$ and on the i.i.d. sequence $(\varepsilon_{t})_{t\in\mathbb{Z}}$ ensuring that a stationary process $(x_{t})_{t\in\mathbb{Z}}$ satisfying $x_{t}=F(x_{t-1},\dots,x_{t-k},\varepsilon_{t})$ exists and is $\beta$-mixing. Lemma C.5 specializes such a result in the context of Gaussian ar(1) processes, which will be considered in the empirical study in Section C.2.
Lemma C.5 (Gaussian ar(1) process).
Consider a generic sequence $(\varepsilon_{t})_{t\in\mathbb{Z}}$ of i.i.d. random variables from a $\mbox{N}(0,\sigma^{2})$. Moreover, let $-1<\theta<1$ and $\psi\in\mathbb{R}$. Then the stationary solution to $x_{t}=\psi+\theta x_{t-1}+\varepsilon_{t}$ is $\beta$-mixing and has coefficients $\beta(k)\leq\smash{|\theta|^{k}/(2\sqrt{1-\theta^{2}})}=(2\sqrt{1-\theta^{2}})^{-1}\exp(-k\log(1/|\theta|))$, $k\in\mathbb{N}$, thus meeting (VII).
Notice that in the empirical study in Section C.2 the focus will be on inference for the ar parameter $\theta$. Clearly, in this case it is not sufficient to focus on the marginal distribution of each $x_{t}$. Rather, one should leverage the bivariate distribution for the pairs $\tilde{x}_{t}:=(x_{t},x_{t+1})$; see also Bernton et al. (2019) where such a strategy is named delay reconstruction. This procedure simply changes the focus to the bivariate stochastic process $(\tilde{x}_{t})_{t\in\mathbb{Z}}$, but does not alter the mixing properties. In particular, if $\beta_{x}(k)$ and $\beta_{\tilde{x}}(k)$ are the mixing coefficients of $(x_{t})_{t\in\mathbb{Z}}$ and $(\tilde{x}_{t})_{t\in\mathbb{Z}}$, respectively, then from Definition C.1 we have $\beta_{\tilde{x}}(k)=\beta_{x}(k-1)$, for $k\geq 1$. Notice that identifiability is a key to ensure concentration in the space of parameters of interest as in Corollary B.1. This motivates further research, beyond the scope of this article, to derive delay reconstruction strategies ensuring identifiability in more complex processes.
Leveraging Lemma C.2 along with the newly-introduced Assumption (VII), Proposition C.6 states convergence of the abc posterior when $\varepsilon_{n}=\varepsilon$ is fixed and $n\to\infty$.
Proposition C.6.
Under Assumptions (III), (IV) and (VII), for any $\varepsilon>\tilde{\varepsilon}$, it holds that
$\pi_{n}^{(\varepsilon)}(\theta)\rightarrow\pi(\theta\mid\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon)\propto\pi(\theta)\mathds{1}\left\{\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon\right\},$ (C.2)
almost surely with respect to $y_{1:n}\sim\mu^{*(n)}$, as $n\rightarrow\infty$.
According to Proposition C.6, replacing Assumption (I) with (VII), does not alter the uniform convergence properties of the abc posterior originally stated in Corollary 3.2 under the i.i.d. assumption. This allows to inherit the discussion after Corollary 3.2 also beyond i.i.d. settings, while suggesting that similar extensions would be possible in the regime $\varepsilon_{n}\rightarrow\varepsilon^{*}$ and $n\rightarrow\infty$. These extensions are stated in Theorem C.7, which provides an important generalization of Theorem 3.3 beyond the i.i.d. case.
Theorem C.7.
Let $\varepsilon_{n}=\varepsilon^{*}+\bar{\varepsilon}_{n}$, and assume (II), (III), (IV) and (VII). Then, if $\bar{\varepsilon}_{n}\to 0$ is such that $\sqrt{n}\bar{\varepsilon}_{n}^{2}\to\infty$ and $\bar{\varepsilon}_{n}/\mathfrak{R}_{s_{n}}(\mathfrak{F})\to\infty$, with $s_{n}=\lfloor n/(2\lfloor\sqrt{n}\rfloor)\rfloor$, we have
$\smash{\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}\Bigl{(}\Bigl{\{}\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})}>\varepsilon^{*}+\frac{4\bar{\varepsilon}_{n}}{3}+2\mathfrak{R}_{s_{n}}(\mathfrak{F})+\frac{4b}{\sqrt{n}}+\Bigl{(}\frac{2b^{2}}{s_{n}}\log\frac{n}{\bar{\varepsilon}_{n}^{L}}\Bigr{)}^{1/2}\Bigr{\}}\Bigr{)}\leq\frac{4\cdot 3^{L}}{c_{\pi}n},$
with $\mathbb{P}_{y_{1:n}}$–probability going to $1$ as $n\rightarrow\infty$, where $\mathfrak{R}_{s_{n}}=\sup_{\mu\in\mathcal{P}(\mathcal{Y})}\mathfrak{R}_{\mu,s_{n}}$.
As for Proposition C.6, also Theorem C.7 shows that informative concentration inequalities similar to those derived in Section 3.2, can be obtained beyond the i.i.d. setting. These results provide insights comparable to those in Theorem 3.3 with the only difference that in this case we require $\sqrt{n}\bar{\varepsilon}_{n}^{2}\to\infty$ rather than $n\bar{\varepsilon}_{n}^{2}\to\infty$ and the term $2b^{2}/n$ within the bound in Theorem 3.3 is now replaced by $2b^{2}/s_{n}$ with $s_{n}\sim\sqrt{n}/2$ as $n\to\infty$. This means that $\bar{\varepsilon}_{n}$ must shrink to zero with a rate at least $n^{1/4}$ slower than the one allowed in the i.i.d. setting. This is an interesting result which clarifies that when moving beyond i.i.d. regimes concentration can still be achieved, although with a slower rate. Such a rate might be pessimistic in some models and we believe it may be improved under future refinements of Lemma C.2.
Notice that (VII) could be relaxed to include $\beta$-mixing processes whose coefficients $\beta(k)$ vanish to zero, but at a non-exponential rate, e.g., $\beta(k)\sim 1/(k+1)^{\xi}$ for some $\xi>0$. In this case, we could still use Lemma C.2 to prove concentration, but with a smaller $s_{n}$, that would lead to even slower rates. However, we did not provide the most general result for the sake of readability. As for processes that are not $\beta$-mixing, we are not aware of results similar to Lemma C.2 in this context. This is an important direction for future research.
C.2 Illustrative simulation in non-i.i.d. settings
Let us illustrate the results in Section C.1 on a simple simulation study focusing on a contaminated Gaussian $\textsc{ar}(1)$ process. More specifically, the uncontaminated data are generated from the model $y^{*}_{t}=0.5y^{*}_{t-1}+\varepsilon_{t}$ for $t=1,\ldots,100$ with $\varepsilon_{t}\sim\mbox{N}(0,1)$ independently, and initial state $y^{*}_{0}\sim\mbox{N}(0,1)$. Then, similarly to the simulation study in Section 5, these data are contaminated with a growing fraction $\alpha\in\{0.05,0.10,0.15\}$ of independent realizations from a $\mbox{N}(20,1)$. As such, each observed data point $y_{t}$ is either equal to $y_{t}^{*}$ or to a sample from $\mbox{N}(20,1)$, for $t=1,\ldots,100$. For Bayesian inference, we assume an $\textsc{ar}(1)$ model $z_{t}=\theta z_{t-1}+\varepsilon_{t}$, with $\varepsilon_{t}\sim\mbox{N}(0,1)$, and focus on learning $\theta$ via discrepancy-based abc under a uniform prior on $[-1,1]$ for $\theta$.
Table C.1: Concentration and runtimes in seconds (for a single discrepancy evaluation) of abc under mmd with Gaussian kernel, Wasserstein-1 distance, summary-based distance (covariance) and kl divergence for an $\textsc{ar}(1)$ Huber contamination model with $\alpha\in\{0.05,0.10,0.15\}$. $\textsc{mse}=\hat{\mathbb{E}}_{\mu^{*(n)}}[\hat{\mathbb{E}}_{\textsc{abc}}(\theta-\theta_{0})^{2}]$, $\theta_{0}=0.5$.
mse ($\alpha=0.05$) mse ($\alpha=0.10$) mse ($\alpha=0.15$) time
(ips) mmd 0.029 0.036 0.049 $<$ 0.01”
(ips) Wasserstein-1 0.043 0.091 0.180 $<$ 0.01”
(ips) summary (covariance) 0.575 0.998 1.001 $<$ 0.01”
(non–ips) kl 0.058 0.060 0.061 $<$ 0.01”
Rejection abc is implemented under the same settings and discrepancies considered in Section 5. However, as discussed in Section C.1, in this case we focus on distances among the empirical distributions of the $n=m=100$ observed $(y_{0},y_{1}),(y_{1},y_{2}),\ldots,(y_{99},y_{100})$ and synthetic $(z_{0},z_{1}),(z_{1},z_{2}),\ldots,(z_{99},z_{100})$ pairs. This is consistent with the delay reconstruction strategy in Bernton et al. (2019) and is motivated by the fact that information on $\theta$ is in the bivariate distributions, rather than in the marginals. For the same reason, in implementing summary-based abc we consider the sample covariance rather than the sample mean.
Table C.1 summarizes the concentration achieved by the different discrepancies analyzed under the aforementioned non-i.i.d. data generating process and model, at varying contamination $\alpha\in\{0.05,0.10,0.15\}$. The results are coherent with those displayed in Table 1 for the i.i.d. scenario and further clarify that discrepancies with guarantees of uniform convergence and concentration generally provide a robust choice, including in non-i.i.d. contexts.
Appendix D Proofs of Theorems, Corollaries and Propositions
Proof of Theorem 3.1.
Note that, by leveraging the first inequality in Lemma 2.6, we have $\mathbb{P}_{y_{1:n}}[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})>2\mathfrak{R}_{\mu^{*},n}(\mathfrak{F})+\delta]\leq\exp(-n\delta^{2}/2b^{2}).$ Hence, setting $\delta=1/n^{1/4}$, and recalling that $\mathfrak{R}_{\mu^{*},n}(\mathfrak{F})\leq\mathfrak{R}_{n}(\mathfrak{F})$, it follows $\mathbb{P}_{y_{1:n}}[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})>2\mathfrak{R}_{n}(\mathfrak{F})+1/n^{1/4}]\leq\exp(-\sqrt{n}/2b^{2})$; note that $\sum_{n\geq 0}\exp(-\sqrt{n}/2b^{2})<\infty$. Therefore, if we define the event
$\mbox{E}_{n}=\{\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})\leq 2\mathfrak{R}_{n}(\mathfrak{F})+1/n^{1/4}\},$ (D.1)
then $\mathds{1}\{\mbox{E}_{n}^{c}\}\rightarrow 0$ almost surely with respect to $y_{1:n}\stackrel{{\scriptstyle\scriptsize\mbox{i.i.d.}}}{{\sim}}\mu^{*}$ as $n\rightarrow\infty$. Now, notice that
$\pi_{n}^{(\varepsilon)}\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon\}\\ =\pi_{n}^{(\varepsilon)}\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon\}\mathds{1}\{\mbox{E}_{n}\}+\pi_{n}^{(\varepsilon)}\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon\}\mathds{1}\{\mbox{E}_{n}^{c}\}.$
Hence, in the following we focus on $\pi_{n}^{(\varepsilon)}\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon\}\mathds{1}\{\mbox{E}_{n}\}$. To this end, recall that
$\pi_{n}^{(\varepsilon)}(\theta)$ $\propto\pi(\theta)\int\mathds{1}\left\{\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\hat{\mu}_{z_{1:n}})\leq\varepsilon\right\}\mu_{\theta}^{n}(dz_{1:n})$
$=\pi(\theta)\int\mathds{1}\left\{\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon+W_{\mathfrak{F}}(z_{1:n})\right\}\mu_{\theta}^{n}(dz_{1:n})=:\pi(\theta)p_{n}(\theta),$
where $W_{\mathfrak{F}}(z_{1:n})=\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})-\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\hat{\mu}_{z_{1:n}})$, whereas $p_{n}(\theta)$ denotes the probability of generating a sample $z_{1:n}$ from $\mu_{\theta}^{n}$ which leads to accept the parameter value $\theta$. Note that, by applying the triangle inequality twice, we have
$-\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})-\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})\leq W_{\mathfrak{F}}(z_{1:n})\leq\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*}),$
and, hence, $|W_{\mathfrak{F}}(z_{1:n})|\leq\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$. This implies that the quantity $p_{n}(\theta)$ can be bounded below and above as follows
$\begin{split}&\int\mathds{1}\left\{\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon-\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})-\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})\right\}\mu_{\theta}^{n}(dz_{1:n})\\ &\qquad\leq p_{n}(\theta)\leq\int\mathds{1}\left\{\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})\right\}\mu_{\theta}^{n}(dz_{1:n}).\end{split}$
Applying again Lemma 2.6 yields $\mathbb{P}_{z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})>2\mathfrak{R}_{\mu_{\theta},n}(\mathfrak{F})+\delta\right]\leq\exp(-n\delta^{2}/2b^{2})$. Therefore, setting $\delta=1/n^{1/4}$, and recalling that $\mathfrak{R}_{\mu_{\theta},n}(\mathfrak{F})\leq\mathfrak{R}_{n}(\mathfrak{F})$ and that we are on the event given in (D.1), it follows
$-\exp(-\sqrt{n}/2b^{2})+\mathds{1}\{\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon-4\mathfrak{R}_{n}(\mathfrak{F})-2/n^{1/4}\}\\ \leq p_{n}(\theta)\leq\exp(-\sqrt{n}/2b^{2})+\mathds{1}\{\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon+4\mathfrak{R}_{n}(\mathfrak{F})+2/n^{1/4}\}.$ (D.2)
Now, notice that the acceptance probability is defined as $p_{n}=\int p_{n}(\theta)\pi(d\theta)$. Hence, integrating with respect to $\pi(\theta)$ in the above inequalities yields, for $n$ large enough,
$\pi\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon-c_{\mathfrak{F}}\}-e_{n}\leq p_{n}\leq\pi\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon+c_{\mathfrak{F}}\}+e_{n},$
where $c_{\mathfrak{F}}=4\lim\sup\mathfrak{R}_{n}(\mathfrak{F})$, as in Equation (3), thus concluding the first part of the proof.
To proceed with the second part of the proof, notice that, by the definition of $\tilde{\varepsilon}=\inf\{\epsilon>0:\smash{\pi\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\epsilon\}}>0\}$, the left part of the above inequality is bounded away from zero for $n$ large enough, whenever $\varepsilon-c_{\mathfrak{F}}>\tilde{\varepsilon}$. This implies that also the acceptance probability $p_{n}$ is strictly positive. As a consequence, for such $n$, it follows that
$\pi_{n}^{(\varepsilon)}(\mbox{A})=\frac{\int p_{n}(\theta)\mathds{1}_{\mbox{\small A}}(\theta)\pi(d\theta)}{\int p_{n}(\theta)\pi(d\theta)}=\frac{\int p_{n}(\theta)\mathds{1}_{\mbox{\small A}}(\theta)\pi(d\theta)}{p_{n}},$
is well-defined for any event A. Then, leveraging the upper bound in (D.2) yields
$\pi_{n}^{(\varepsilon)}\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})>\varepsilon+4\mathfrak{R}_{n}(\mathfrak{F})\}=\frac{\int p_{n}(\theta)\mathds{1}\left\{\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})>\varepsilon+4\mathfrak{R}_{n}(\mathfrak{F})\right\}\pi(d\theta)}{\int p_{n}(\theta)\pi(d\theta)}$
$\qquad\leq\frac{\int\mathds{1}\left\{\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon+4\mathfrak{R}_{n}(\mathfrak{F})+2/n^{1/4}\right\}\mathds{1}\left\{\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})>\varepsilon+4\mathfrak{R}_{n}(\mathfrak{F})\right\}\pi(d\theta)}{p_{n}}$
$\qquad\qquad+\frac{\int\exp(-\sqrt{n}/2b^{2})\mathds{1}\left\{\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})>\varepsilon+4\mathfrak{R}_{n}(\mathfrak{F})\right\}\pi(d\theta)}{p_{n}}.$
To conclude the proof it is now necessary to control both terms. Note that we already proved that the denominator $p_{n}$ is bounded away from zero for $n$ large enough. Both numerators are bounded by $1$, and going to 00 when $n\rightarrow\infty$. Thus, by the dominated convergence theorem, both summands in the above upper bound for $\pi_{n}^{(\varepsilon)}\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})>\varepsilon+4\mathfrak{R}_{n}(\mathfrak{F})\}$ go to zero. This implies, as a direct consequence, that $\smash{\pi_{n}^{(\varepsilon)}}\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon+4\mathfrak{R}_{n}(\mathfrak{F})\}\rightarrow 1$, almost surely with respect to $y_{1:n}\stackrel{{\scriptstyle\scriptsize\mbox{i.i.d.}}}{{\sim}}\mu^{*}$ as $n\rightarrow\infty$, thereby concluding the proof. ∎
Proof of Corollary 3.2.
Note that by combining Equation (2) in Lemma 2.6 with the result $\sum_{n>0}\exp[-n\delta^{2}/(2b^{2})]<\infty$, the Borel–Cantelli Lemma implies that both $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})$ and $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$ converge to 0 almost surely when $\mathfrak{R}_{n}(\mathfrak{F})\rightarrow 0$ as $n\rightarrow\infty$. Hence, since
[…]
it follows that $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\rightarrow\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ almost surely as $n\rightarrow\infty$.
Combining this result with the proof of Theorem 1 in Jiang, Wu and Wong (2018) yields the statement of Corollary 3.2. Notice that, as discussed in Section 3.1, the limiting pseudo-posterior in Corollary 3.2 is well-defined only for those $\varepsilon>\tilde{\varepsilon}$, with $\tilde{\varepsilon}$ as in Theorem 3.1. ∎
Proof of Theorem 3.3.
Since Lemma 2.6 and $\mathfrak{R}_{n}(\mathfrak{F})=\sup_{\mu\in\mathcal{P}(\mathcal{Y})}\mathfrak{R}_{\mu,n}(\mathfrak{F})\geq\mathfrak{R}_{\mu,n}(\mathfrak{F})$ hold for every $\mu\in\mathcal{P}(\mathcal{Y})$, then, for every integer $n\geq 1$ and any scalar $\delta\geq 0$, Equation (2) implies $\mathbb{P}_{x_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\leq 2\mathfrak{R}_{n}(\mathfrak{F})+\delta\right]\geq 1-\exp(-n\delta^{2}/2b^{2}).$ Moreover, since this result holds for any $\delta\geq 0$, it follows that $\mathbb{P}_{x_{1:n}}[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\leq 2\mathfrak{R}_{n}(\mathfrak{F})+(c_{1}-2\mathfrak{R}_{n}(\mathfrak{F}))]\geq 1-\exp[-n(c_{1}-2\mathfrak{R}_{n}(\mathfrak{F}))^{2}/2b^{2}],$ for any $c_{1}\geq 2\mathfrak{R}_{n}(\mathfrak{F})$. Hence,
$\mathbb{P}_{x_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\leq c_{1}\right]\geq 1-\exp[-n(c_{1}-2\mathfrak{R}_{n}(\mathfrak{F}))^{2}/2b^{2}].$ (D.3)
Recalling the settings of Theorem 3.3, consider the sequence $\bar{\varepsilon}_{n}\to 0$ as $n\to\infty$, with $n\bar{\varepsilon}^{2}_{n}\to\infty$ and $\bar{\varepsilon}_{n}/\mathfrak{R}_{n}(\mathfrak{F})\to\infty$, which is possible by Assumption (IV). These regimes imply that $\bar{\varepsilon}_{n}$ goes to zero slower than $\mathfrak{R}_{n}(\mathfrak{F})$ and, hence, for $n$ large enough, $\bar{\varepsilon}_{n}/3>2\mathfrak{R}_{n}(\mathfrak{F})$. Therefore, under Assumptions (I)–(III) it is now possible to apply (D.3) to $y_{1:n}$, by setting $c_{1}=\bar{\varepsilon}_{n}/3$, which yields
$\mathbb{P}_{y_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})\leq\bar{\varepsilon}_{n}/3\right]\geq 1-\exp[-n(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{n}(\mathfrak{F}))^{2}/2b^{2}].$
Since $-n(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{n}(\mathfrak{F}))^{2}=-n\bar{\varepsilon}^{2}_{n}[1/9+4(\mathfrak{R}_{n}(\mathfrak{F})/\bar{\varepsilon}_{n})^{2}-(4/3)\mathfrak{R}_{n}(\mathfrak{F})/\bar{\varepsilon}_{n}]$, it follows that $-n(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{n}(\mathfrak{F}))^{2}\rightarrow-\infty$ when $n\to\infty$. From the above settings we also have that $n\bar{\varepsilon}^{2}_{n}\to~{}\infty$ and $\mathfrak{R}_{n}(\mathfrak{F})/\bar{\varepsilon}_{n}\to 0$, when $n\to\infty$. Therefore, as a consequence, we obtain $1-\exp[-n(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{n}(\mathfrak{F}))^{2}/2b^{2}]\rightarrow 1$ as $n\to\infty$. Hence, in the rest of this proof, we will restrict to the event $\{\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})\leq\bar{\varepsilon}_{n}/3\}$.
Denote with $\mathbb{P}_{\theta,z_{1:n}}$ the joint distribution of $\theta\sim\pi$ and $z_{1:n}$ i.i.d. from $\mu_{\theta}$. By definition of conditional probability, for any $c_{2}$, including $c_{2}>2\mathfrak{R}_{n}(\mathfrak{F})$, it follows that
$\begin{split}&\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}\left(\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})>\varepsilon^{*}+4\bar{\varepsilon}_{n}/3+c_{2}\}\right)\\ &\qquad=\frac{\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})>\varepsilon^{*}+4\bar{\varepsilon}_{n}/3+c_{2},\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}\right]}{\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}\right]}.\end{split}$ (D.4)
To derive an upper bound for the above ratio, we first identify an upper bound for its numerator. In addressing this goal, we leverage the triangle inequality $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$, since $\mathcal{D}_{\mathfrak{F}}$ is a semimetric, and the previously-proved result that the event $\{\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})\leq\bar{\varepsilon}_{n}/3\}$ has $\mathbb{P}_{y_{1:n}}$–probability going to $1$, thereby obtaining
$\mathbb{P}_{\theta,z_{1:n}}[\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})>\varepsilon^{*}+4\bar{\varepsilon}_{n}/3+c_{2},\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}]$
$\qquad\quad\leq\mathbb{P}_{\theta,z_{1:n}}[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})>\varepsilon^{*}+4\bar{\varepsilon}_{n}/3+c_{2},$
$\qquad\quad\qquad\qquad\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}]$
$\qquad\quad\leq\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})>\bar{\varepsilon}_{n}/3+c_{2}\right]\leq\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})>c_{2}\right].$
Rewriting $\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})>c_{2}\right]$ as $\int_{\theta\in\Theta}\mathbb{P}_{z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})>c_{2}\mid\theta\right]\pi({\rm d}\theta)$ and applying (D.3) to $z_{1:n}$ yields
$\begin{split}{\int_{\theta\in\Theta}}&{\mathbb{P}_{z_{1:n}}}[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}}{,}\mu_{\theta})>c_{2}{\mid}\theta]\pi({\rm d}\theta)=\int_{\theta\in\Theta}(1-{\mathbb{P}_{z_{1:n}}}[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}}{,}\mu_{\theta})\leq c_{2}{\mid}\theta])\pi({\rm d}\theta)\\ &\quad\leq\int_{\theta\in\Theta}\exp[-n(c_{2}-2\mathfrak{R}_{n}(\mathfrak{F}))^{2}/2b^{2}]\pi({\rm d}\theta)=\exp[-n(c_{2}-2\mathfrak{R}_{n}(\mathfrak{F}))^{2}/2b^{2}].\end{split}$
Hence, the numerator of the ratio in Equation (D.4) can be upper bounded by $\exp[-n(c_{2}-2\mathfrak{R}_{n}(\mathfrak{F}))^{2}/2b^{2}]$ for any $c_{2}>2\mathfrak{R}_{n}(\mathfrak{F})$. As for the denominator, defining the event $\mbox{E}_{n}:=\{\theta\in\Theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}/3\}$ and applying again the triangle inequality, we have that
$\mathbb{P}_{\theta,z_{1:n}}[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon^{*}{+}\ \bar{\varepsilon}_{n}]\geq\int_{{\footnotesize\mbox{E}}_{n}}\mathbb{P}_{z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon^{*}{+}\ \bar{\varepsilon}_{n}\mid\theta\right]\pi({\rm d}\theta)$
$\qquad\geq\int_{{\footnotesize\mbox{E}}_{n}}\mathbb{P}_{z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})+\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}\mid\theta\right]\pi({\rm d}\theta)$
$\qquad\geq\int_{{\footnotesize\mbox{E}}_{n}}\mathbb{P}_{z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})\leq\bar{\varepsilon}_{n}/3\mid\theta\right]\pi({\rm d}\theta),$
where the last inequality follows directly from the fact that it is possible to restrict to the event $\{\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})\leq\bar{\varepsilon}_{n}/3\}$, and that we are integrating over $\mbox{E}_{n}:=\{\theta\in\Theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}/3\}$. Applying again (D.3) to $z_{1:n}$, with $c_{1}=\bar{\varepsilon}_{n}/3>2\mathfrak{R}_{n}(\mathfrak{F})$, the last term of the above inequality can be further lower bounded by
${\int_{{\footnotesize\mbox{E}}_{n}}}{(}1-\exp[-n(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{n}(\mathfrak{F}))^{2}/2b^{2}])\pi({\rm d}\theta)$
$\qquad\qquad\qquad=\pi(\mbox{E}_{n})(1-\exp[-n(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{n}(\mathfrak{F}))^{2}/2b^{2}]),$
with $\pi(\mbox{E}_{n})\geq c_{\pi}(\bar{\varepsilon}_{n}/3)^{L}$ by (II), and, as shown before, $1-\exp[-n(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{n}(\mathfrak{F}))^{2}/2b^{2}]\to 1$, when $n\to\infty$, which implies, for $n$ large enough, $21-\exp[-n(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{n}(\mathfrak{F}))^{2}/2b^{2}]>1/$ 2. Leveraging both results, the denominator in (D.4) is lower bounded by $(c_{\pi}/2)\left(\bar{\varepsilon}_{n}/{3}\right)^{L}$. Let us now combine the upper and lower bounds derived, respectively, for the numerator and the denominator of the ratio in (D.4), to obtain
[…]
with $\mathbb{P}_{y_{1:n}}$–probability going to $1$ as $n\rightarrow\infty$. To conclude the proof it suffices to replace $c_{2}$ in (D.5) with $2\mathfrak{R}_{n}(\mathfrak{F})+\sqrt{(2b^{2}/n)\log(M_{n}/\bar{\varepsilon}_{n}^{L})}$, which is never lower than $2\mathfrak{R}_{n}(\mathfrak{F})$. Finally, setting $M_{n}=n$ yields the statement of Theorem 3.3. ∎
Proof of Corollary B.1.
Corollary B.1 follows by replacing the bounds in the proof of Corollary 1 by Bernton et al. (2019) with the newly-derived ones in Theorem 3.3. ∎
Proof of Corollary 4.1.
Recall that in the case of mmd with kernels bounded by $1$ we have $\mathfrak{R}_{n}(\mathfrak{F})\leq n^{-1/2}$. Hence, regarding the upper and lower bounds on $p_{n}$ in (3) it holds
$\begin{split}&\pi\{\theta:\mathcal{D}_{\textsc{mmd}}(\mu_{\theta},\mu^{*})\leq\varepsilon-c_{\mathfrak{F}}\}\geq\pi\{\theta:\mathcal{D}_{\textsc{mmd}}(\mu_{\theta},\mu^{*})\leq\varepsilon-4/\sqrt{n}\},\\ &\pi\{\theta:\mathcal{D}_{\textsc{mmd}}(\mu_{\theta},\mu^{*})\leq\varepsilon+c_{\mathfrak{F}}\}\leq\pi\{\theta:\mathcal{D}_{\textsc{mmd}}(\mu_{\theta},\mu^{*})\leq\varepsilon+4/\sqrt{n}\}.\end{split}$
Combining the above inequalities with the result in (3), and taking the limit for $n\rightarrow\infty$, proves the first part of the statement. The second part is a direct application of Corollary 3.2 to the case of mmd with bounded kernels, after noticing that the aforementioned inequality $\mathfrak{R}_{n}(\mathfrak{F})\leq n^{-1/2}$ implies $\mathfrak{R}_{n}(\mathfrak{F})\rightarrow 0$ as $n\rightarrow\infty$. ∎
Proof of Corollary 4.2.
To prove Corollary 4.2, it suffices to plug $\bar{\varepsilon}_{n}=[(\log n)/n]^{\frac{1}{2}}$ and ${b=1}$ into the statement of Theorem 3.3, and then upper-bound the resulting radius via the inequalities $\mathfrak{R}_{n}(\mathfrak{F})\leq n^{-1/2}$ and $\log n\geq 1$. The latter holds for any $n\geq 3$ and hence for $n\rightarrow\infty$. ∎
Proof of Proposition 4.3.
We first show that, under $(A1)$$(A3)$, Assumptions 1 and 2 made in Bernton et al. (2019) are satisfied under mmd when $f_{n}(\bar{\varepsilon}_{n})=1/(n\bar{\varepsilon}^{2}_{n})$ and $c(\theta)=\mathbb{E}_{z}\left[k(z,z)\right]$, with $z\sim\mu_{\theta}$.
Consistent with the above goal, first recall that, by standard properties of mmd,
$\mathcal{D}^{2}_{\textsc{mmd}}(\mu_{1},\mu_{2})=\mathbb{E}_{x_{1},x_{1}^{\prime}}\left[k(x_{1},x_{1}^{\prime})\right]-2\mathbb{E}_{x_{1},x_{2}}\left[k(x_{1},x_{2})\right]+\mathbb{E}_{x_{2},x_{2}^{\prime}}\left[k(x_{2},x_{2}^{\prime})\right],$ (D.6)
with $x_{1},x_{1}^{\prime}\sim\mu_{1}$ and $x_{2},x_{2}^{\prime}\sim\mu_{2}$, all independently; see e.g., Chérief-Abdellatif and Alquier (2022). Since $k(x,x^{\prime})=\left<\phi(x),\phi(x^{\prime})\right>_{\mathcal{H}}$ (see e.g., Muandet et al., 2017), the above result implies that
$\begin{split}&\mathcal{D}^{2}_{\textsc{mmd}}(\mu_{1},\mu_{2})\\ &=\mathbb{E}_{x_{1},x_{1}^{\prime}}\left[\left<\phi(x_{1}),\phi(x_{1}^{\prime})\right>_{\mathcal{H}}\right]-2\mathbb{E}_{x_{1},x_{2}}\left[\left<\phi(x_{1}),\phi(x_{2})\right>_{\mathcal{H}}\right]+\mathbb{E}_{x_{2},x_{2}^{\prime}}\left[\left<\phi(x_{2}),\phi(x_{2}^{\prime})\right>_{\mathcal{H}}\right]\\ &=||\mathbb{E}_{x_{1}}[\phi(x_{1})]||_{\mathcal{H}}^{2}-2\left<\mathbb{E}_{x_{1}}[\phi(x_{1})],\mathbb{E}_{x_{2}}[\phi(x_{2})]\right>_{\mathcal{H}}+||\mathbb{E}_{x_{2}}[\phi(x_{2})]||_{\mathcal{H}}^{2}\\ &=\left\|\mathbb{E}_{x_{1}}\left[\phi(x_{1})\right]-\mathbb{E}_{x_{2}}\left[\phi(x_{2})\right]\right\|^{2}_{\mathcal{H}}.\end{split}$ (D.7)
Leveraging Equations (D.6)–(D.7) and basic Markov inequalities, for any $\bar{\varepsilon}_{n}\geq 0$, it holds
$\mathbb{P}_{y_{1:n}}\left[\mathcal{D}_{\textsc{mmd}}(\hat{\mu}_{y_{1:n}},\mu^{*})>\bar{\varepsilon}_{n}\right]\leq(1/\bar{\varepsilon}_{n}^{2})\mathbb{E}_{y_{1:n}}\left[\mathcal{D}^{2}_{\textsc{mmd}}(\hat{\mu}_{y_{1:n}},\mu^{*})\right]$
$=(1/\bar{\varepsilon}_{n}^{2})\mathbb{E}_{y_{1:n}}[\left\|(1/n){\textstyle{\sum\nolimits_{i=1}^{n}}}\phi(y_{i})-\mathbb{E}_{y}\left[\phi(y)\right]\right\|_{\mathcal{H}}^{2}]\leq[1/(n^{2}\bar{\varepsilon}_{n}^{2})]{\textstyle{\sum\nolimits_{i=1}^{n}}}\mathbb{E}_{y_{i}}[\left\|\phi(y_{i})\right\|_{\mathcal{H}}^{2}]$
$\leq[1/(n\bar{\varepsilon}_{n}^{2})]\mathbb{E}_{y_{1}}[\left\|\phi(y_{1})\right\|_{\mathcal{H}}^{2}]=[1/(n\bar{\varepsilon}_{n}^{2})]\mathbb{E}_{y_{1}}\left[k(y_{1},y_{1})\right]=[1/(n\bar{\varepsilon}_{n}^{2})]\mathbb{E}_{y}\left[k(y,y)\right],$
with $y\sim\mu^{*}$. Since $[1/(n\bar{\varepsilon}_{n}^{2})]\mathbb{E}_{y}\left[k(y,y)\right]\rightarrow 0$ as $n\rightarrow\infty$ by condition $(A1)$, we have that $\mathcal{D}_{\textsc{mmd}}(\hat{\mu}_{y_{1:n}},\mu^{*})\to 0$ in $\mathbb{P}_{y_{1:n}}$–probability as $n\to\infty$, thus meeting Assumption 1 in Bernton et al. (2019). Moreover, as a direct consequence of the above derivations,
$\mathbb{P}_{z_{1:n}}\left[\mathcal{D}_{\textsc{mmd}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})>\bar{\varepsilon}_{n}\right]\leq[1/(n\bar{\varepsilon}_{n}^{2})]\mathbb{E}_{z}\left[k(z,z)\right].$
Thus, setting $1/(n\bar{\varepsilon}^{2}_{n})=f_{n}(\bar{\varepsilon}_{n})$ and $\mathbb{E}_{z}\left[k(z,z)\right]=c(\theta)$, with $z\sim\mu_{\theta}$, ensures that
$\mathbb{P}_{z_{1:n}}\left[\mathcal{D}_{\textsc{mmd}}(\hat{\mu}_{z_{1:n}},\mu_{\theta}]>\bar{\varepsilon}_{n}\right]\leq c(\theta)f_{n}(\bar{\varepsilon}_{n}),$
with $f_{n}(u)=1/(nu^{2})$ strictly decreasing in $u$ for any fixed $n$, and $f_{n}(u)\rightarrow 0$ as $n\rightarrow\infty$, for fixed $u$. Moreover, by Assumptions $(A2)$$(A3)$, $c(\theta)=\mathbb{E}_{z}\left[k(z,z)\right]$ is $\pi$-integrable and there exist a $\delta_{0}>0$ and a $c_{0}>0$ such that $c(\theta)<c_{0}$ for any $\theta$ satisfying $(\mathbb{E}_{z,z^{\prime}}\left[k(z,z^{\prime})\right]-2\mathbb{E}_{z,y}\left[k(y,z)\right]+\mathbb{E}_{y,y^{\prime}}\left[k(y,y^{\prime})\right])^{1/2}=\mathcal{D}_{\textsc{mmd}}(\mu_{\theta},\mu^{*})\leq\varepsilon^{*}+\delta_{0}$. This ensures that Assumption 2 in Bernton et al. (2019) holds.
Finally, note that Assumption 3 in Bernton et al. (2019), is verified by our Assumption (II). Therefore, under the assumptions in Proposition 4.3 it is possible to apply Proposition 3 in Bernton et al. (2019) with $f_{n}(\bar{\varepsilon}_{n})=1/(n\bar{\varepsilon}^{2}_{n})$, $c(\theta)=\mathbb{E}_{z}\left[k(z,z)\right]$ and $R=M_{n}$, which yields the concentration result in Proposition 4.3. ∎
Proof of Proposition 4.4.
Assumptions (A1’) and (A2’) imply that the norm $\|\|x\|\|_{\psi_{1}}$ (see Equation (14) in Lei (2020) for a definition) is uniformly bounded when $x\sim\mu_{\theta}$ and $x\sim\mu^{*}$. Thus, we can use (15) in Lei (2020) to obtain
$\mathbb{P}_{z_{1:n}}\left(\mathcal{D}_{\textrm{wass}}(\mu_{\theta},\hat{\mu}_{z_{1:n}})>u\right)\leq\exp[-c^{\prime}n(u-c_{1}n^{-1/\max(d,3)})_{+}^{2}]=:f_{n}(u)$
and, similarly, $\mathbb{P}_{y_{1:n}}\left(\mathcal{D}_{\textrm{wass}}(\mu^{*},\hat{\mu}_{y_{1:n}})>u\right)\leq f_{n}(u)$, where $c^{\prime}$ depends only on the constant $c$ in (A1’) and (A2’), and $c_{1}$ is a universal constant. Notice that (15) in Lei (2020) requires $d>2$. If $d=1$, we can define $x^{\prime}=(x,0,0)\in\mathbb{R}^{3}$ and apply the result in $\mathbb{R}^{3}$ (we can proceed similarly if $d=2$). This is why Proposition 4.4 is stated with $\max(d,3)$. The above bounds ensure that Assumptions 1–2 in Bernton et al. (2019) are met. Moreover, condition (II) verifies Assumption 3 in Bernton et al. (2019). Thus, we can apply Proposition 3 of Bernton et al. (2019) with $f_{n}(\cdot)$ defined as above and under the vanishing conditions on $\bar{\varepsilon}_{n}$ in Proposition 4.4. This implies that when $n\rightarrow\infty$, and $\bar{\varepsilon}_{n}\rightarrow 0$ such that $f_{n}(\bar{\varepsilon}_{n})\rightarrow 0$, for some ${C\in(0,\infty)}$ and any $M_{n}\in(0,\infty)$, with $\mathbb{P}_{y_{1:n}}$–probability going to $1$ as $n\rightarrow\infty$, it holds
$\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}(\{\theta:\mathcal{D}_{\textrm{wass}}(\mu_{\theta},\mu^{*})>\varepsilon^{*}+4\bar{\varepsilon}_{n}/3+f_{n}^{-1}\left(\bar{\varepsilon}_{n}^{L}/M_{n}\right)\})\leq C/M_{n}.$
Recall that our restriction $n^{-1/\max(d,3)}\ll\bar{\varepsilon}_{n}$, together with $n\bar{\varepsilon}^{2}_{n}\to\infty$, imply that for $n$ large enough $f_{n}(\cdot)$ is invertible and $f_{n}(\bar{\varepsilon}_{n})\rightarrow 0$. On such a range, we have that $f_{n}^{-1}\left(\bar{\varepsilon}_{n}^{L}/M_{n}\right)=[(1/(c^{\prime}n))\log(M_{n}/\bar{\varepsilon}_{n}^{L})]^{1/2}+c_{1}n^{-{1}/{\max(d,3)}},$ which concludes the proof. ∎
Proof of Lemma C.2.
Let $(x_{t})_{t\in\mathbb{Z}}$ be a stochastic process with $\beta$-mixing coefficients $\beta(k)$, for $k\in\mathbb{N}$. Moreover, denote with $\mu^{(n)}$ and $\mu$ the joint distribution of a sample $x_{1:n}$ from $(x_{t})_{t\in\mathbb{Z}}$ and its constant marginal $\mu=\mu^{(1)}$, respectively. Then, by combining Proposition 2 and Lemma 2 in Mohri and Rostamizadeh (2008) under our notation, we have that for any $b$-uniformly bounded class $\mathfrak{F}$, any integer $n\geq 1$ and any scalar $\delta\geq 0$, the inequality
$\mathbb{P}_{x_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)>2\mathfrak{R}_{\mu,n/(2K)}(\mathfrak{F})+\delta\right]\\ \leq 2\exp(-n\delta^{2}/Kb^{2})+2(n/2K-1)\beta(K),$
holds for every integer $K>0$ such that $n/(2K)\in\mathbb{N}$, where $\mathfrak{R}_{\mu,n/(2K)}$ is the Rademacher complexity based an i.i.d. sample of size $n/(2K)$ from $\mu$; see Definition 2.5. The above concentration inequality also implies
$\begin{split}&\mathbb{P}_{x_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\leq 2\mathfrak{R}_{\mu,n/(2K)}(\mathfrak{F})+\delta\right]\\ &\qquad\qquad\qquad\qquad\qquad\geq 1-2\exp(-n\delta^{2}/Kb^{2})-2(n/2K-1)\beta(K)\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\geq 1-2\exp[-n\delta^{2}/(4Kb^{2})]-2(n/2K)\beta(K).\end{split}$ (D.8)
Notice that, in order to ensure that both $2\exp(-n\delta^{2}/4Kb^{2})$ and $2(n/2K)\beta(K)$ vanish to zero (under Assumption (VII) for $\beta(K)$), it is tempting to apply (D.8) with $K=n^{\alpha}$ for some $10<\alpha<$ 1. Unfortunately, there is no reason for such a $K$ to be an integer. A solution, would be to let $K=\lfloor n^{\alpha}\rfloor$, but in this case $n/(2K)$ might not be an integer. To address such issues, it is necessary to consider a careful modification of (D.8). To this end write the Euclidean division $n=n^{\prime}+r$ where $n^{\prime}=2Kh\leq n$, $h=\lfloor n/(2K)\rfloor$ and $0\leq r<2K$. Then, under the common marginal assumption and recalling also Proposition 1 in Mohri and Rostamizadeh (2008) together with the triangle inequality and the fact that the functions $f$ within $\mathfrak{F}$ are $b$-uniformly bounded, we have that
$\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)$ $=\sup_{f\in\mathfrak{F}}\left|\frac{1}{n}\sum\nolimits_{i=1}^{n}[f(x_{i})-\mathbb{E}_{\mu}f(x)]\right|$
$\leq\sup_{f\in\mathfrak{F}}\left|\frac{1}{n}\sum\nolimits_{i=1}^{n^{\prime}}[f(x_{i})-\mathbb{E}_{\mu}f(x)]\right|+\frac{1}{n}\sup_{f\in\mathfrak{F}}\left|\sum\nolimits_{i=n^{\prime}+1}^{n^{\prime}+r}[f(x_{i})-\mathbb{E}_{\mu}f(x)]\right|$
$\leq\sup_{f\in\mathfrak{F}}\left|\frac{1}{n^{\prime}}\sum\nolimits_{i=1}^{n^{\prime}}[f(x_{i})-\mathbb{E}_{\mu}f(x)]\right|+\frac{2br}{n}\leq\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n^{\prime}}},\mu)+\frac{4bK}{n},$
where the last inequality follows directly from the definition of $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n^{\prime}}},\mu)$ together with the fact that $0\leq r<2K$. Applying now (D.8) to $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n^{\prime}}},\mu)$ yields
$\begin{split}&\mathbb{P}_{x_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n^{\prime}}},\mu)+4bK/n\leq 2\mathfrak{R}_{\mu,h}(\mathfrak{F})+\delta+4bK/n\right]\geq 1-2\exp(-h\delta^{2}/2b^{2})-2h\beta(K).\end{split}$
Therefore, since $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\leq\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n^{\prime}}},\mu)+4bK/n$ we also have that
$\begin{split}&\mathbb{P}_{x_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\leq 2\mathfrak{R}_{\mu,h}(\mathfrak{F})+\delta+4bK/n\right]\geq 1-2\exp(-h\delta^{2}/2b^{2})-2h\beta(K).\end{split}$
To conclude the proof, notice that to prove convergence and concentration of the abc posterior it will be sufficient to let $K=\lfloor\sqrt{n}\rfloor$. Therefore, by replacing $K=\lfloor\sqrt{n}\rfloor$ in the above inequality and within the expression for $h=\lfloor n/(2K)\rfloor$ we have
$\mathbb{P}_{x_{1:n}}\Bigl{[}\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\leq 2\mathfrak{R}_{\mu,s_{n}}(\mathfrak{F})+\frac{4b}{\sqrt{n}}+\delta\Bigr{]}\geq 1-2\exp(-s_{n}\delta^{2}/2b^{2})-2s_{n}\beta(\lfloor\sqrt{n}\rfloor)$
where $s_{n}=\lfloor n/(2\lfloor\sqrt{n}\rfloor)\rfloor$ and the term $4b/\sqrt{n}$ follows directly from the fact that $\lfloor\sqrt{n}\rfloor/n\leq\sqrt{n}/n=1/\sqrt{n}$. ∎
Proof of Lemma C.5.
The proof follows the arguments used in Chapter 2 of Doukhan (1994) to study general Markov chains. In particular, when $(x_{t})_{t\in\mathbb{Z}}$ is a stationary Markov chain with invariant distribution $\pi(\cdot)$ and transition kernel $P(\cdot,\cdot)$, a result proven in Davydov (1974) and recalled in page 87–88 of Doukhan (1994) gives $\beta(k)=\mathbb{E}_{x\sim\pi}\|P^{k}(x,\cdot)-\pi(\cdot)\|_{\textsc{TV}}.$ When $-1<\theta<1$, standard results for the ar(1) model in Lemma C.5 lead to the invariant distribution $\pi=\mbox{N}(\psi/(1-\theta),\sigma^{2}/(1-\theta^{2}))$.
As for $P^{k}(x,\cdot)$, notice that, under such an ar(1) model with starting point $x$, we can write
$x_{k}$ $=\theta x_{k-1}+\psi+\varepsilon_{k}=\theta(\theta x_{k-2}+\psi+\varepsilon_{k-1})+\psi+\varepsilon_{k}=\dots$
$=\theta^{k}x+\psi\sum\nolimits_{l=0}^{k-1}\theta^{l}+\sum\nolimits_{l=0}^{k-1}\theta^{l}\varepsilon_{k-l}.$
Therefore, $P^{k}(x,\cdot)=\mbox{N}(\theta^{k}x+\psi\sum_{l=0}^{k-1}\theta^{l},\sigma^{2}\sum_{l=0}^{k-1}\theta^{2l})$.
Moreover, notice that, by direct application of standard properties of finite power series, we have $\sum_{l=0}^{k-1}\theta^{l}=(1-\theta^{k})/(1-\theta)$ and $\sum_{l=0}^{k-1}\theta^{2l}=\smash{(1-\theta^{2k})/(1-\theta^{2})}$. Since our goal is to derive an upper bound for $\beta(k)$ and provided that the kl divergence among Gaussian densities is available in closed form, let us first consider the Pinsker’s inequality
$\|P^{k}(x,\cdot)-\pi(\cdot)\|_{\textsc{TV}}\leq[\mathcal{D}_{\textrm{KL}}(P^{k}(x,\cdot),\pi(\cdot))/2]^{1/2}$
where $\mathcal{D}_{\textrm{KL}}$ stands for the kl divergence. Since both $P^{k}(x,\cdot)$ and $\pi(\cdot)$ are Gaussian, then
$\mathcal{D}_{\textrm{KL}}(P^{k}(x,\cdot),\pi(\cdot))$
$=\frac{1}{2}\left[\frac{\sigma^{2}(1-\theta^{2k})}{\sigma^{2}}-1+\frac{[\theta^{k}x+\psi(1-\theta^{k})/(1-\theta)-\psi/(1-\theta)]^{2}}{\sigma^{2}/(1-\theta^{2})}+\log\frac{\sigma^{2}}{\sigma^{2}(1-\theta^{2k})}\right]$
$=\frac{1}{2}\left[(1-\theta^{2k})-1+\frac{\theta^{2k}[x-\psi/(1-\theta)]^{2}}{\sigma^{2}/(1-\theta^{2})}+\log\frac{1}{1-\theta^{2k}}\right]$
$=\frac{1}{2}\left[-\theta^{2k}+\frac{\theta^{2k}[x-\psi/(1-\theta)]^{2}}{\sigma^{2}/(1-\theta^{2})}+\log\left(1+\frac{\theta^{2k}}{1-\theta^{2k}}\right)\right]$
$\leq\frac{1}{2}\left[-\theta^{2k}+\frac{\theta^{2k}[x-\psi/(1-\theta)]^{2}}{\sigma^{2}/(1-\theta^{2})}+\frac{\theta^{2k}}{1-\theta^{2k}}\right]\leq\frac{\theta^{2k}}{2}\left[-1+\frac{[x-\psi/(1-\theta)]^{2}}{\sigma^{2}/(1-\theta^{2})}+\frac{1}{1-\theta^{2}}\right].$
Therefore, by leveraging the above result, together with standard properties of the expectation, we have
$\beta(k)$ $\leq\mathbb{E}_{x\sim\pi}[\mathcal{D}_{\textrm{KL}}(P^{k}(x,\cdot),\pi(\cdot))/2]^{1/2}\leq[\mathbb{E}_{x\sim\pi}\mathcal{D}_{\textrm{KL}}(P^{k}(x,\cdot),\pi(\cdot))/2]^{1/2}$
$\leq\left[\frac{\theta^{2k}}{4}\left[-1+\frac{\mathbb{E}_{x\sim\pi}[x-\psi/(1-\theta)]^{2}}{\sigma^{2}/(1-\theta^{2})}+\frac{1}{1-\theta^{2}}\right]\right]^{1/2}=\sqrt{\frac{\theta^{2k}}{4(1-\theta^{2})}}=\frac{|\theta|^{k}}{2\sqrt{1-\theta^{2}}},$
which concludes the proof. ∎
Proof of Proposition C.6.
Under Assumption (VII), for any fixed $\delta>0$, we have that $\sum\nolimits_{n>0}[2\exp(-s_{n}\delta^{2}/2b^{2})+2s_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi})]<\infty.$ Therefore, combining Lemma C.2 with Assumption (IV), both $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})$ and $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$ converge to 0 almost surely as $n\to\infty$, by the Borel–Cantelli Lemma. As a result, since
[…]
it holds that $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\rightarrow\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ almost surely as $n\rightarrow\infty$. To conclude it suffices to apply again the proof of Theorem 1 in Jiang, Wu and Wong (2018); see also the proof of Corollary 3.2. ∎
Proof of Theorem C.7.
To prove Theorem C.7 we will follow the same line of reasoning as in the proof of Theorem 3.3. However, in this case we leverage Lemma C.2 instead of Lemma 2.6. To this end, letting $\delta=c_{1}-2\mathfrak{R}_{s_{n}}(\mathfrak{F})-4b/\sqrt{n}$ with $c_{1}\geq 2\mathfrak{R}_{s_{n}}(\mathfrak{F})+4b/\sqrt{n}$ and $\mathfrak{R}_{s_{n}}=\sup_{\mu\in\mathcal{P}(\mathcal{Y})}\mathfrak{R}_{\mu,s_{n}}$, we obtain, under Assumption (VII), Equation (D.9) below, instead of (D.3).
$\mathbb{P}_{x_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{x_{1:n}},\mu)\leq c_{1}\right]\\ \geq 1-2\exp[-s_{n}(c_{1}-2\mathfrak{R}_{s_{n}}(\mathfrak{F})-4b/\sqrt{n})^{2}/2b^{2}]-2s_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi}).$ (D.9)
As in Theorem 3.3, let $c_{1}=\bar{\varepsilon}_{n}/3$ and notice that, by the settings of Theorem C.7, for $n$ large enough $\bar{\varepsilon}_{n}/3>2\mathfrak{R}_{s_{n}}(\mathfrak{F})+4b/\sqrt{n}$. Therefore, applying (D.9) to $y_{1:n}$, with $c_{1}=\bar{\varepsilon}_{n}/3$, leads to the following upper bound
$\mathbb{P}_{y_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})\leq\bar{\varepsilon}_{n}/3\right]\\ \geq 1-2\exp[-s_{n}(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{s_{n}}(\mathfrak{F})-4b/\sqrt{n})^{2}/2b^{2}]-2s_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi}).$
Recall that, under the settings of Theorem C.7, we have that $\sqrt{n}\bar{\varepsilon}^{2}_{n}\to\infty$ and $\bar{\varepsilon}_{n}/\mathfrak{R}_{s_{n}}(\mathfrak{F})\to\infty$ and, therefore,
$s_{n}(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{s_{n}}(\mathfrak{F})-4b/\sqrt{n})^{2}\sim s_{n}\bar{\varepsilon}^{2}_{n}/9\sim\sqrt{n}\bar{\varepsilon}^{2}_{n}\to\infty.$
Combining this result with the fact that $2s_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi})\to 0$ as $n\to\infty$, it follows that the lower bound for $\mathbb{P}_{y_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})\leq\bar{\varepsilon}_{n}/3\right]$ goes to $1$ as $n\to\infty$. Hence, in the remaining part of the proof, we will restrict to the event $\{\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})\leq\bar{\varepsilon}_{n}/3\}$.
Let $\mathbb{P}_{\theta,z_{1:n}}$ corresponds to the joint distribution of $\theta\sim\pi$ and $z_{1:n}$ from $\mu_{\theta}^{(n)}$. Then, as a direct consequence of the definition of conditional probability, for every positive $c_{2}$, including $c_{2}>2\mathfrak{R}_{s_{n}}(\mathfrak{F})+4b/\sqrt{n}$, it follows that
$\begin{split}&\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}\left(\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})>\varepsilon^{*}+4\bar{\varepsilon}_{n}/3+c_{2}\}\right)\\ &\qquad=\frac{\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})>\varepsilon^{*}+4\bar{\varepsilon}_{n}/3+c_{2},\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}\right]}{\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}\right]}.\end{split}$ (D.10)
To upper bound the ratio in (D.10), let us first derive an upper bound for the numerator. To this end, consider the triangle inequality $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$ (recall that $\mathcal{D}_{\mathfrak{F}}$ is a semimetric), along with the previously-proved result that the event $\{\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})\leq\bar{\varepsilon}_{n}/3\}$ has $\mathbb{P}_{y_{1:n}}$–probability going to $1$. Hence, for $n$ large enough we have
$\mathbb{P}_{\theta,z_{1:n}}[\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})>\varepsilon^{*}+4\bar{\varepsilon}_{n}/3+c_{2},\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}]$
$\qquad\leq\mathbb{P}_{\theta,z_{1:n}}[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})>\varepsilon^{*}+4\bar{\varepsilon}_{n}/3+c_{2},$
$\qquad\qquad\qquad\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}]$
$\qquad\leq\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})>\bar{\varepsilon}_{n}/3+c_{2}\right]\leq\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})>c_{2}\right],$
where $\mathbb{P}_{\theta,z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})>c_{2}\right]=\int_{\theta\in\Theta}\mathbb{P}_{z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})>c_{2}\mid\theta\right]\pi({\rm d}\theta).$
Therefore, leveraging the above result and applying (D.9) to $z_{1:n}$ yields,
$\begin{split}{\int_{\theta\in\Theta}}&{\mathbb{P}_{z_{1:n}}}[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})>c_{2}\mid\theta]\pi({\rm d}\theta)=\int_{\theta\in\Theta}(1-{\mathbb{P}_{z_{1:n}}}[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})\leq c_{2}\mid\theta])\pi({\rm d}\theta)\\ &\qquad\qquad\leq 2\exp[-s_{n}(c_{2}-2\mathfrak{R}_{s_{n}}(\mathfrak{F})-4b/\sqrt{n})^{2}/2b^{2}]+2s_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi}).\end{split}$
This controls the numerator in (D.10). As for the denominator, defining the event $\mbox{E}_{n}:=\{\theta\in\Theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}/3\}$ and applying again the triangle inequality, we have that
$\mathbb{P}_{\theta,z_{1:n}}[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon^{*}{+}\ \bar{\varepsilon}_{n}]\geq\int_{{\footnotesize\mbox{E}}_{n}}\mathbb{P}_{z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\leq\varepsilon^{*}{+}\ \bar{\varepsilon}_{n}\mid\theta\right]\pi({\rm d}\theta)$
$\qquad\geq\int_{{\footnotesize\mbox{E}}_{n}}\mathbb{P}_{z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})+\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})+\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}\mid\theta\right]\pi({\rm d}\theta)$
$\qquad\geq\int_{{\footnotesize\mbox{E}}_{n}}\mathbb{P}_{z_{1:n}}\left[\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})\leq\bar{\varepsilon}_{n}/3\mid\theta\right]\pi({\rm d}\theta).$
The last inequality follows from that fact that we can restrict to the event $\{\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})\leq\bar{\varepsilon}_{n}/3\}$, and that we are integrating over $\mbox{E}_{n}:=\{\theta\in\Theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})\leq\varepsilon^{*}+\bar{\varepsilon}_{n}/3\}$. Let us now apply again (D.9) to $z_{1:n}$, with $c_{1}=\bar{\varepsilon}_{n}/3$, to further lower bound the last term of the above inequality by
${\int_{{\footnotesize\mbox{E}}_{n}}}[1-2\exp[-s_{n}(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{s_{n}}(\mathfrak{F})-4b/\sqrt{n})^{2}/2b^{2}]-2s_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi})]\pi({\rm d}\theta)$
$=\pi(\mbox{E}_{n})[1-2\exp[-s_{n}(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{s_{n}}(\mathfrak{F})-4b/\sqrt{n})^{2}/2b^{2}]-2s_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi})].$
Note that, by Assumption (II), $\pi(\mbox{E}_{n})\geq c_{\pi}(\bar{\varepsilon}_{n}/3)^{L}$. Moreover, as shown before, the quantity $1-2\exp[-s_{n}(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{s_{n}}(\mathfrak{F})-4b/\sqrt{n})^{2}/2b^{2}]-2s_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi})$ goes to $1$ when $n\to\infty$, which also implies that, for a large enough $n$,
$21-2\exp[-s_{n}(\bar{\varepsilon}_{n}/3-2\mathfrak{R}_{s_{n}}(\mathfrak{F})-4b/\sqrt{n})^{2}/2b^{2}]-2s_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi})>1/$ 2 .
Therefore, leveraging both results, the denominator in (D.10) can be lower bounded by the term $(c_{\pi}/2)\left(\bar{\varepsilon}_{n}/{3}\right)^{L}$.
To proceed with the proof, let us combine the upper and lower bounds derived, respectively, for the numerator and the denominator of the ratio in (D.10). This yields, for any integer $K$,
$\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}(\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})>\varepsilon^{*}+4\bar{\varepsilon}_{n}/3+c_{2}\})\\ \leq\frac{2\exp[-s_{n}(c_{2}-2\mathfrak{R}_{s_{n}}(\mathfrak{F})-4b/\sqrt{n})^{2}/2b^{2}]+2s_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi})}{(c_{\pi}/2)(\bar{\varepsilon}_{n}/3)^{L}},$ (D.11)
with $\mathbb{P}_{y_{1:n}}$–probability going to $1$ as $n\rightarrow\infty$.
Replacing $c_{2}$ in (D.11) with $2\mathfrak{R}_{s_{n}}(\mathfrak{F})+\sqrt{(2b^{2}/s_{n})\log(n/\bar{\varepsilon}_{n}^{L})}+4b/\sqrt{n}$ , gives
$\smash{\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}\Bigl{(}\Bigl{\{}\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})}>\smash{\varepsilon^{*}+\frac{4\bar{\varepsilon}_{n}}{3}+2\mathfrak{R}_{s_{n}}(\mathfrak{F})+\frac{4b}{\sqrt{n}}+\Bigl{(}\frac{2b^{2}}{s_{n}}\log\frac{n}{\bar{\varepsilon}_{n}^{L}}\Bigr{)}^{1/2}}\Bigr{\}}\Bigr{)}\\ \leq\frac{4\cdot 3^{L}}{nc_{\pi}}\Bigl{(}1+\frac{ns_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi})}{\bar{\varepsilon}_{n}^{L}}\Bigr{)}.$
To conclude, note that
$ns_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi})/\bar{\varepsilon}_{n}^{L}=n^{1+L/2}s_{n}C_{\beta}\exp(-\gamma\lfloor\sqrt{n}\rfloor^{\xi})/((n\bar{\varepsilon}_{n}^{2})^{L/2}),$
where the numerator goes to 00 and the denominator goes to $\infty$ when $n\to\infty$, under the setting of Theorem C.7. Therefore, with $\mathbb{P}_{y_{1:n}}$–probability going to $1$ as $n\to\infty$, we have
$\smash{\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}\Bigl{(}\Bigl{\{}\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})}>\smash{\varepsilon^{*}+\frac{4\bar{\varepsilon}_{n}}{3}+2\mathfrak{R}_{s_{n}}(\mathfrak{F})+\frac{4b}{\sqrt{n}}+\Bigl{(}\frac{2b^{2}}{s_{n}}\log\frac{n}{\bar{\varepsilon}_{n}^{L}}\Bigr{)}^{1/2}}\Bigr{\}}\Bigr{)}\\ \leq\frac{4\cdot 3^{L}}{nc_{\pi}},$
concluding the proof. ∎